### v1

# Global analysis of charm mixing parameters and determination of the CKM angle $\mathbf{\gamma}$

Abstract
We present an updated global analysis of beauty decays sensitive to the angle $\gamma$ of the Cabibbo-Kobayashi-Maskawa matrix and of $D$-meson mixing data in the framework of approximate universality for charm mixing. We extract the fundamental theoretical parameters determining absorptive and dispersive contributions to $D$ meson mixing and CP violation, together with the angle $\gamma$. The results for the charm mixing parameters are $x_{12}\simeq x=(0.402\pm 0.034)\%$ and $y_{12}\simeq y=(0.629\pm 0.021)\%$, while the two CP-violating phases are given by $\phi_{2}^{M}=(2.4\pm 1.6)\degree$ and $\phi_{2}^{\Gamma}=(2.6\pm 1.2)\degree$. The angle $\gamma$ is found to be $\gamma=(68.4\pm 2.6)\degree$.
I Introduction
Charm mixing occurs through Flavour Changing Neutral Currents (FCNC) that are absent at the tree-level in the Standard Model (SM) and suppressed by the Glashow-Iliopoulous-Maiani (GIM) mechanism and by the hierarchical structure of the Cabibbo-Kobayashi-Maskawa (CKM) matrix. Thus, heavy new physics (NP) may appear in the neutral $D$ system Grossman et al. (2007); Ciuchini et al. (2007); Golowich et al. (2007); Fajfer et al. (2007); Bona et al. (2008); Artuso et al. (2008); Brod et al. (2012a); Pirtskhalava and Uttayarat (2012); Bhattacharya et al. (2012); Cheng and Chiang (2012); Franco et al. (2012); Brod et al. (2012b); Giudice et al. (2012); Feldmann et al. (2012); Keren-Zur et al. (2013); Grossman and Schacht (2019); Dery and Nir (2019); Buras et al. (2021); Schacht and Soni (2022); Schacht (2023) through observables as the amount of CP violation. The latter could be used as an interesting benchmark of the SM since it is suppressed by the fourth power of the sine of the Cabibbo angle $\theta_{C}$.
The experimental study of neutral $D$ mixing has been continuously progressing in the past two decades Bitenc et al. (2008); Aubert et al. (2009); del Amo Sanchez et al. (2010); Asner et al. (2012); Insler et al. (2012); Aaltonen et al. (2013); Peng et al. (2014); Malde et al. (2015); Aaij et al. (2016a, b, c, 2019a); Betti (2021); Ablikim et al. (2021); Aaij et al. (2021a); Ablikim et al. (2024); Pajero (2022); Zhang et al. (2006); Aubert et al. (2007); Aaij et al. (2017, 2024a); Aitala et al. (1999); Link et al. (2000); Csorna et al. (2002); Zupanc et al. (2009); Lees et al. (2013); Ablikim et al. (2015); Starič et al. (2016); Aaij et al. (2019b); Nayak et al. (2020); Aaij et al. (2022, 2023a); Ablikim et al. (2022); Aaij et al. (2019c, 2021b, 2023b, a, 2024a). The most precise estimates of the charm parameters are obtained from the rates of Cabibbo Favoured (CF) (i.e. proportional to $V_{cs}V_{ud}^{*}$) or Doubly Cabibbo Suppressed (DCS) (i.e. proportional to $V_{cd}V_{us}^{*}$) $D$ decays Zhang et al. (2006); Aubert et al. (2007); Aaij et al. (2017, 2024a), from charm decays to CP eigenstates Aitala et al. (1999); Link et al. (2000); Csorna et al. (2002); Zupanc et al. (2009); Lees et al. (2013); Ablikim et al. (2015); Starič et al. (2016); Aaij et al. (2019b); Nayak et al. (2020); Aaij et al. (2022, 2023a); Ablikim et al. (2022) and from Dalitz plot analyses of the three-body mode $K^{0}_{S}\pi^{+}\pi^{-}$ Aaij et al. (2019c, 2021b, 2023b). In particular, LHCb has achieved impressive results during the LHC Run II, including the observation of direct CP asymmetries in the Singly Cabibbo Suppressed (SCS) decays of $D$ mesons to $K^{+}K^{-}$ and $\pi^{+}\pi^{-}$ Aaij et al. (2019a, 2023a) and the very recent measurements of CP-violating observables in $D^{0}\to K^{\pm}\pi^{\mp}$ Aaij et al. (2024a).
This remarkable increase in the experimental accuracy calls for an update of our previous works on $D$ meson mixing Bevan et al. (2012, 2014), generalizing the so-called “superweak assumption”, in which the weak phase describing CP violation in the interference between $D$ meson decays with and without absorptive mixing is set to zero, to the more suitable “approximate universality” scenario introduced in ref. Kagan and Silvestrini (2021) to be detailed below. To fully exploit the large amount of experimental information we include also beauty observables in the combination, allowing for a simultaneous determination of the CKM angle $\gamma=\arg(-V_{ud}V_{ub}^{*}V_{cb}V_{cd}^{*})$ from tree-level $B$ decays Abe et al. (2006); Horii et al. (2011); Nayak et al. (2013); Aaij et al. (2015a, b, 2016d); LHC (2016); Aaij et al. (2016e, 2019d, 2021c, 2020); Adachi et al. (2023, 2024a); Aaij et al. (2024b, c); Poluektov et al. (2010); Aaij et al. (2016f); Resmi et al. (2019); Aaij et al. (2021d); Abudinén et al. (2022); Aaij et al. (2023c, 2024d, d, 2024e, 2018, 2021e, 2024f) and of $D$ mixing parameters. Our results can be compared with the latest LHCb Aaij et al. (2021f); LHC (2024) estimates, the Belle/Belle II analysis Adachi et al. (2024b) for $\gamma$ and HFLAV Amhis et al. (2023) for the charm parameters. The consistency of the fit is checked by comparing $\gamma$ estimates obtained by combining measurements involving different types of $B$ mesons separately: charged $B^{\pm}$ or neutral $B^{0}$ and $B^{0}_{s}$.
The paper is organized as follows: the dispersive-absorptive formalism for $D$ meson mixing in the framework of approximate universality is briefly summarized in Sec. II. The expressions of the charm observables are described in Sec. III, while the beauty observables are introduced in Sec. IV and their dependence on the parameters of interest of this study are reported in Apps. A and B. We present and discuss the most relevant results in Sec. V. A complete list of the parameters introduced and of their prior distributions is reported in App. C, while their posterior probability intervals are given in App. D. A summary can be found in Sec. VI.
II Charm Mixing Formalism
In this section, we present a parametrization of the quantities needed to describe charm decays and $D$ mixing underlying all the observables considered in the combination.
We indicate with $M$ ($\overline{M}$) a generic $B$ ($\overline{B}$) or $D$ ($\overline{D}$) meson, while its decay amplitudes to CP conjugate final states $f$ and $\overline{f}$ are denoted as
$\mathcal{A}_{M}^{f}=\bra{f}\mathcal{H}\ket{M},$ $\quad\quad\mathcal{A}^{\overline{f}}_{M}=\bra{\overline{f}}\mathcal{H}\ket{M},$ (1)
$\mathcal{A}^{f}_{\overline{M}}=\bra{f}\mathcal{H}\ket{\overline{M}},$ $\quad\quad\mathcal{A}^{\overline{f}}_{\overline{M}}=\bra{\overline{f}}\mathcal{H}\ket{\overline{M}},$
where $\mathcal{H}$ is the relevant effective Hamiltonian involved in the process.
The time evolution of a linear combination of $D^{0}$ and $\overline{D^{0}}$ mesons follows the Schrödinger equation, with a $2\times 2$ non-Hermitian Hamiltonian $\mathbf{H}$ (see e.g. Nir (1992); Branco et al. (1999); Grossman et al. (2009); Kagan and Sokoloff (2009)) that in terms of its dispersive $\mathbf{M}$ and absorptive $\mathbf{\Gamma}$ components is given by
$\mathbf{H}=\mathbf{M}-i/2\mathbf{\Gamma},$ (2)
so that its matrix element between charm states can be written as
$\langle D^{0}|\mathbf{H}|\overline{D^{0}}\rangle=H_{12}=M_{12}-i/2\Gamma_{12}.$ (3)
Adopting the long-lived (L) and short-lived (S) meson notation, the Hamiltonian eigenstates read
|$\ket{D_{\text{L,S}}}=p|D^{0}\rangle\pm q|\overline{D^{0}}\rangle,$ (4)
with $p$ and $q$ complex coefficients satisfying $\left|p\right|^{2}+\left|q\right|^{2}=1$. Then, the differences between masses and decay widths of the Hamiltonian eigenstates are parametrized through the usual quantities
$x=\frac{m_{\text{S}}-m_{\text{L}}}{\Gamma},\quad\quad\quad\quad\quad y=\frac{\Gamma_{\text{S}}-\Gamma_{\text{L}}}{2\Gamma},$ (5)
with $\Gamma$ being the averaged $D^{0}$ lifetime.
$D$ meson mixing is described in terms of the matrix elements of $\mathbf{H}$ by the parameters
$\phi_{12}=\arg\bigg{(}\frac{M_{12}}{\Gamma_{12}}\bigg{)},\qquad x_{12}=2\frac{|M_{12}|}{\Gamma}\qquad\mathrm{and}\qquad y_{12}=\frac{|\Gamma_{12}|}{\Gamma}.$ (6)
The phase $\phi_{12}$ governs CP violation in pure mixing, while $x_{12}$ and $y_{12}$ are CP-conserving.
The CP-violating quantity $|q/p|-1$ and the observables in Eq. (5) are related to the mixing parameters in Eq. (6) as
$|x|=x_{12},\qquad\qquad\left|\frac{q}{p}\right|-1=\frac{x_{12}y_{12}}{x_{12}^{2}+y_{12}^{2}}\sin\phi_{12},\qquad\qquad|y|=y_{12},$ (7)
up to negligible corrections quadratic in $\sin\phi_{12}$.
In the SM, it is possible to decompose the dispersive and absorptive parts of $H_{12}$ in terms of U-spin amplitudes as follows Kagan and Silvestrini (2021):
$\xi_{12}^{\text{SM}}=\frac{(\lambda^{s}_{uc}-\lambda^{d}_{uc})^{2}}{4}\xi_{2}+\frac{(\lambda^{s}_{uc}-\lambda^{d}_{uc})\lambda^{b}_{uc}}{2}\xi_{1}+\frac{(\lambda^{b}_{uc})^{2}}{4}\xi_{0},\qquad\xi=M,\Gamma,$ (8)
[…]
In the approximate universality scenario, CP violation in the interference between decays with and without dispersive and absorptive mixing can be described through two final state-independent weak phases $\phi_{2}^{M,\Gamma}$, defined as
$\phi_{2}^{\xi}=\arg\frac{\xi_{12}}{\frac{1}{4}(\lambda^{s}_{uc}-\lambda^{d}_{uc})^{2}\xi_{2}},\qquad\xi=M,\Gamma.$ (9)
Furthermore, the difference between $\phi_{2}^{M}$ and $\phi_{2}^{\Gamma}$ is exactly $\phi_{12}$, meaning that all types of indirect CP violation concerning $D$ meson mixing, namely CPV in mixing and in the interference between mixing and decay, are entirely determined by knowledge of these two universal weak phases.
Employing in the definition of Eq. (9) the expressions of $M_{12}^{\text{SM}}$ and $\Gamma_{12}^{\text{SM}}$ given in Eq. (8) and neglecting the third (smallest) contribution, the orders of magnitude of the two CP-violating phases in the SM can be roughly estimated to be Kagan and Silvestrini (2021)
$(\phi_{2}^{M})\sim(\phi_{2}^{\Gamma})\simeq\left|\frac{\lambda^{b}_{uc}}{\theta_{C}}\right|\sin\gamma\times\frac{\Gamma_{1}}{\Gamma_{2}}\sim\left|\frac{\lambda^{b}_{uc}}{\theta_{C}}\right|\sin\gamma\times\frac{1}{\varepsilon}\sim(2.2\times 10^{-3})\times\bigg{(}\frac{0.3}{\varepsilon}\bigg{)}.$ (10)
From Eqs. (6) and (8), we get the relation $\left|\Gamma_{2}\right|\simeq y_{12}\ \Gamma/\left|\lambda^{s}_{uc}\right|^{2}$, which can be employed in Eq. (10) to provide an upper bound on $\left|\phi_{2}^{\Gamma}\right|$ as Kagan and Silvestrini (2021)
$\left|\phi_{2}^{\Gamma}\right|\simeq\frac{\left|\lambda^{b}_{uc}\lambda^{s}_{uc}\sin\gamma\right|}{y_{12}}\times\frac{\left|\Gamma_{1}\right|}{\Gamma}<(5\times 10^{-3})\times\bigg{(}\frac{0.63\%}{y_{12}}\bigg{)},$ (11)
where we used the conservative assumption $\left|4\Gamma_{1}/(\Gamma_{0}-\Gamma_{2})\right|<1$, a ratio which is nominally of $\mathcal{O}(\varepsilon)$.
For a given final state $f$, CP violation in the interference between decays with and without mixing can be equivalently described by the weak phase of the observable
$\lambda^{f}_{D}\equiv\frac{q}{p}\frac{\mathcal{A}_{\overline{D}}^{f}}{\mathcal{A}_{D}^{f}},$ (12)
which becomes universal in the framework of approximate universality and we denote by $\phi_{2}$. The posterior of $\phi_{2}$ is determined from the charm parameters as
$\tan 2\phi_{2}=-\bigg{(}\frac{x_{12}^{2}\sin 2\phi_{2}^{M}+y_{12}^{2}\sin 2\phi_{2}^{\Gamma}}{x_{12}^{2}\cos 2\phi_{2}^{M}+y_{12}^{2}\cos 2\phi_{2}^{\Gamma}}\bigg{)}.$ (13)
Notice that $\phi_{2}$ corresponds to the phase $\phi$ extracted in global analyses under the assumption of final state independence (see e.g. refs. Aaij et al. (2021f); LHC (2024)).
In the superweak assumption, dispersive mixing is supposed to be the dominant source of CPV in the charm system, corresponding to the approximation where $\phi_{2}^{\Gamma}=0$ and $\phi_{12}=\phi_{2}^{M}$ or, equivalently, $\phi=\phi_{2}=\arctan\frac{1-|q/p|^{2}}{1+|q/p|^{2}}\frac{x}{y}$. Then, going from the superweak assumption, adopted in our previous works Bevan et al. (2012, 2014), to approximate universality, used in this work, involves moving from a description in terms of a single independent CPV parameter ($\phi_{12}$ or $\phi_{2}$) to a scenario with two free CPV parameters ($\phi_{2}^{M}$ and $\phi_{2}^{\Gamma}$ or $\phi_{2}$ and $|q/p|-1$).
III Charm observables
The charm observables used in our combination rely on time-integrated and time-dependent studies of $D$ meson decays, reconstructed from two- and multi-body final states. The final states can be CP eigenstates or originated from CF or DCS decays. We parametrize the decay amplitudes through the ratios of their magnitudes and strong phases and provide expressions for the observables in Tab. 1 in terms of charm decay, mixing and CP-violating parameters in the framework of approximate universality for the neutral $D$ system.
III.1 CF/DCS decays to $\mathbf{K^{\pm}X}$
In this section, we consider CF/DCS decays to $K^{\pm}X$, such as $D^{0}\to K^{\pm}\pi^{\mp}$. We denote with $f$ the final state of the CF decay of the $D^{0}$ meson (e.g. $K^{-}\pi^{+}$) and by “wrong-sign” (WS) and “right-sign” (RS) the time-dependent DCS (i.e. $D^{0}\to\overline{f}$ and $\overline{D^{0}}\to f$) and CF (i.e. $D^{0}\to f$ and $\overline{D^{0}}\to\overline{f}$) decays respectively. We parametrize the ratios of decay amplitudes in Eq. (1) as
$r_{D}^{f}e^{-i\delta_{D}^{f}}=\frac{\mathcal{A}^{\overline{f}}_{D}}{\mathcal{A}^{f}_{D}}=\frac{\mathcal{A}^{f}_{\overline{D}}}{\mathcal{A}^{\overline{f}}_{\overline{D}}},$ (14)
where we have neglected possible new weak phases in CF/DCS decays (weak phases are fully negligible in the SM).
The observables used in the fit for this type of final states are extracted from the ratios of the WS and RS decay rates Aaij et al. (2017); Aubert et al. (2007); Zhang et al. (2006), given by
$\frac{\Gamma(D^{0}\to\overline{f},t)}{\Gamma(D^{0}\to f,t)}=$ $\ (r_{D}^{f})^{2}+r_{D}^{f}y_{f}^{\prime+}\tau+\frac{(x_{f}^{\prime+})^{2}+(y_{f}^{\prime+})^{2}}{4}\tau^{2},$ (15)
$\frac{\Gamma(\overline{D^{0}}\to f,t)}{\Gamma(\overline{D^{0}}\to\overline{f},t)}=$ $\ (r_{D}^{f})^{2}+r_{D}^{f}y_{f}^{\prime-}\tau+\frac{(x_{f}^{\prime-})^{2}+(y_{f}^{\prime-})^{2}}{4}\tau^{2},$
up to second order in the charm mixing and CP-violating parameters and with $\tau=\Gamma t$. The coefficients $x_{f}^{\prime\pm}$ and $y_{f}^{\prime\pm}$ can be written as a rotation of an angle $\phi_{2}$ of their CP-conserving limits, which we denote with $x^{\prime}_{f}$ and $y^{\prime}_{f}$, respectively, as
$\begin{pmatrix}x_{f}^{\prime\pm}\\ y_{f}^{\prime\pm}\end{pmatrix}=-\left|\frac{q}{p}\right|^{\pm 1}\begin{pmatrix}\cos\phi_{2}&\pm\sin\phi_{2}\\ \mp\sin\phi_{2}&\cos\phi_{2}\end{pmatrix}\begin{pmatrix}x^{\prime}_{f}\\ y^{\prime}_{f}\end{pmatrix},$ (16)
with
$x_{f}^{\prime}=$ $\ x\cos\delta_{D}^{f}+y\sin\delta_{D}^{f},$ (17)
$y_{f}^{\prime}=$ $\ y\cos\delta_{D}^{f}-x\sin\delta_{D}^{f}.$
The linear and quadratic terms in Eq. (15) can be decomposed into
Observables $\mathbf{D^{0}}$ decays Ref. Observables $\mathbf{D^{0}}$ decays Ref.
$x^{\prime\pm}_{K\pi\pi^{0}}$, $y^{\prime\pm}_{K\pi\pi^{0}}$ $D\to K\pi\pi^{0}$ Aubert et al. (2009) $(x^{2}+y^{2})/2$ $D\to Kl\overline{\nu}_{l}$ Bitenc et al. (2008)
$(r_{D}^{K\pi})^{2}$, $x^{2}$, $y$, $\frac{\mathcal{B}(D^{0}\to K^{0}_{s}K^{+}\pi^{-})}{\mathcal{B}(D^{0}\to K^{0}_{s}K^{-}\pi^{+})}$,
$\cos\delta_{D}^{K\pi}$, $D\to K\pi$ Asner et al. (2012) $\delta_{D}^{K^{0}_{S}K\pi}$, $D\to K^{0}_{S}K\pi$ Insler et al. (2012)
$\sin\delta_{D}^{K\pi}$ $\kappa_{D}^{K^{0}_{S}K\pi}$
$\tilde{y}_{\text{CP}}$ $D\to XX$ Aitala et al. (1999); Link et al. (2000); Csorna et al. (2002); Zupanc et al. (2009); Lees et al. (2013); Ablikim et al. (2015); Starič et al. (2016); Aaij et al. (2019b); Nayak et al. (2020); Aaij et al. (2022) $x$, $y$, $|q/p|$, $\phi_{2}$ $D\to K^{0}_{S}\pi\pi$ Peng et al. (2014)
$F_{+}^{XX\pi^{0}}$ $D\to XX\pi^{0}$ Malde et al. (2015) $x$, $y$ $D\to K^{0}_{S}\pi\pi$ del Amo Sanchez et al. (2010); Aaij et al. (2016a)
$F_{+}^{4\pi}$ $D\to 4\pi$ Malde et al. (2015); Ablikim et al. (2024) $\frac{\mathcal{B}(D\to K^{0}_{s}K^{+}\pi^{-})}{\mathcal{B}(D^{0}\to K^{0}_{s}K^{-}\pi^{+})}$ $D\to K^{0}_{S}K\pi$ Aaij et al. (2016b)
$(r_{D}^{K\pi})^{2}$, $(x^{\prime\pm}_{K\pi})^{2}$, $y^{\prime\pm}_{K\pi}$ $D\to K\pi$ Zhang et al. (2006); Aubert et al. (2007); Aaij et al. (2017) $(x^{2}+y^{2})/4$ $D\to K3\pi$ Aaij et al. (2016c)
$x_{\text{CP}}$, $y_{\text{CP}}$, $\Delta x$, $\Delta y$ $D\to K^{0}_{S}\pi\pi$ Aaij et al. (2019c, 2021b, 2023b) $\Delta A_{D}^{\text{CP}}$, $(\left\langle\tau\right\rangle^{KK}_{E^{\prime}}\pm\left\langle\tau\right\rangle^{\pi\pi}_{E^{\prime}})$ $D\to XX$ Aaij et al. (2019a)
$r_{D}^{K3\pi}$, $\delta_{D}^{K3\pi}$, $\kappa_{D}^{K3\pi}$, $D\to K3\pi$ Ablikim et al. (2021) $\Delta Y^{KK}-\Delta Y^{\pi\pi}$ $D\to XX$ Aaij et al. (2021a)
$r_{D}^{K\pi\pi^{0}}$, $\delta_{D}^{K\pi\pi^{0}}$, $\kappa_{D}^{K\pi\pi^{0}}$ $D\to K\pi\pi^{0}$ $\Delta Y$
$(r_{D}^{K\pi})^{2}$, $c^{(^{\prime})}_{K\pi}$, $\Delta c^{(^{\prime})}_{K\pi}$, $D\to K\pi$ Aaij et al. (2024a) $A_{D}(f^{+})$, $A_{D}(\pi\pi\pi^{0})$, $D\to f^{+}$, $D\to K\pi$ Ablikim et al. (2022)
$\Delta\tilde{c}^{(^{\prime})}_{K\pi}$ $r_{D}^{K\pi}\cos\delta_{D}^{K\pi}$, $r_{D}^{K\pi}\sin\delta_{D}^{K\pi}$ $D\to XX\pi^{0}$
$A_{D}^{\text{CP}}(KK)$, $\left\langle\tau\right\rangle_{E}^{KK}$ $D\to K^{+}K^{-}$ Aaij et al. (2023a)
Table 1: Charm observables used in the combination. We use $XX$ to indicate both $\pi^{+}\pi^{-}$ and $K^{+}K^{-}$ states, while $K\pi$ stands for $K^{\mp}\pi^{\pm}$. We refer to the multi-body final states $K^{\mp}\pi^{\pm}\pi^{+}\pi^{-}$ and $\pi^{+}\pi^{-}\pi^{+}\pi^{-}$ as $K3\pi$ and $4\pi$, respectively. The state $f^{+}$ stands for a CP-even final state. A description of the observables can be found in Secs. III.1, III.2 and III.3.
their CP-conserving ($c_{f}$, $c^{\prime}_{f}$) and CP-violating parts ($\Delta c_{f}$, $\Delta c^{\prime}_{f}$), as follows:
$y_{f}^{\prime\pm}=c_{f}\pm$ $\ \Delta c_{f},$ (18)
$\frac{(x_{f}^{\prime\pm})^{2}+(y_{f}^{\prime\pm})^{2}}{4}=$ $\ c_{f}^{\prime}\pm\Delta c_{f}^{\prime},$
where the coefficients on the right-hand side of Eq. (18) are given in terms of the charm mixing parameters as
$c_{f}=x_{12}\cos\phi_{2}^{M}\sin\delta_{D}^{f}-y_{12}\cos\phi_{2}^{\Gamma}\cos\delta_{D}^{f},$ $\qquad\Delta c_{f}=-x_{12}\sin\phi_{2}^{M}\cos\delta_{D}^{f}-y_{12}\sin\phi_{2}^{\Gamma}\sin\delta_{D}^{f},$ (19)
$c^{\prime}_{f}=\frac{x_{12}^{2}+y_{12}^{2}}{4}+\frac{(r_{D}^{f})^{2}}{4}(y_{12}^{2}-x_{12}^{2}),$ $\qquad\Delta c^{\prime}_{f}=\frac{1}{2}x_{12}y_{12}\sin\phi_{12}.$
The separation of the CP-violating from the CP-conserving contributions entering the WS/RS ratios of $D^{0}\to K^{\pm}\pi^{\mp}$ decays has been performed for the first time this year by LHCb Aaij et al. (2024a), combining both the Run 1 and Run 2 results. However, while the Run 1 measurements can be used directly to extract the coefficients in Eq. (19), the Run 2 data use experimental information from $D^{0}\to K^{+}K^{-}$ decays to correct a detection asymmetry. This operation leaves invariant $c_{f}$ and $c^{\prime}_{f}$, while it changes the expressions of the CPV observables as
$\Delta\tilde{c}_{f}=$ $\ \Delta c_{f}-c_{f}a_{D}^{KK}-2r_{D}^{f}\Delta Y^{KK},$ (20)
$\Delta\tilde{c}^{\prime}_{f}=$ $\ \Delta c^{\prime}_{f}-2c^{\prime}_{f}a_{D}^{KK}-2r_{D}^{f}c_{f}\Delta Y^{KK},$
and the parameters $r_{D}^{f}$ are replaced by $r_{D}^{f}(1\mp a_{D}^{KK})$ in Eq. (15). Here, $a_{D}^{KK}$ is the direct CP asymmetry in $D^{0}\to K^{+}K^{-}$ decays and we introduced the parameter $\Delta Y^{KK}$. They were measured in refs. Aaij et al. (2021a) and Aaij et al. (2023a), and are defined in terms of decay amplitudes and decay rates in Eqs. (23) and (26), respectively, when describing the observables involving SCS decays to CP eigenstates.
We consider in the combination also observables coming from the analysis of quantum correlated $D^{0}-\overline{D^{0}}$ pairs (see e.g. Ablikim et al. (2022)). Here, one of the $D$ mesons decays to the so-called tagging mode, while the other decays to the signal mode, typically $K^{-}\pi^{+}$. Then, the following asymmetries are measured:
$A_{D}(f^{+})=\frac{\mathcal{B}(D_{f^{+}}\to K^{-}\pi^{+})-\mathcal{B}(D_{f^{-}}\to K^{-}\pi^{+})}{\mathcal{B}(D_{f^{+}}\to K^{-}\pi^{+})+\mathcal{B}(D_{f^{-}}\to K^{-}\pi^{+})},$ (21)
where $\mathcal{B}(D_{f^{\pm}}\to K^{-}\pi^{+})$ are the branching fractions of neutral charm states decays to $K^{-}\pi^{+}$ when tagged by CP-even or multi-body modes, also known as quasi-CP eigenstates, $D_{f^{+}}$ or CP-odd modes $D_{f^{-}}$. For example, we could have $f^{+}=\pi^{+}\pi^{-}(\pi^{0})$ and $f^{-}=K^{0}_{S}\pi^{0}$. In the CP-conserving limit, the asymmetry in Eq. (21) depends on the charm parameters as
$A_{D}(f^{+})=\frac{F_{+}^{f^{+}}(y-2r_{D}^{K\pi}\cos\delta_{D}^{K\pi})}{1+(r_{D}^{K\pi})^{2}+(1-F_{+}^{f^{+}})(y+2r_{D}^{K\pi}\cos\delta_{D}^{K\pi})},$ (22)
up to $\mathcal{O}(x,y,(r_{D}^{K\pi})^{2})$ corrections. Here, we have introduced the so-called CP-even fraction $F_{+}^{f^{+}}$ of the mode $f^{+}$, which is a non-negative real number needed for quasi-CP eigenstates that reaches its maximum of one for CP-even eigenstates. A definition of $F_{+}^{f^{+}}$ in terms of decay amplitudes is given in Sec. III.3, in the context of multi-body final states.
III.2 SCS decays to CP eigenstates
SCS decays to CP eigenstates $f$, such as $D^{0}\to\pi^{+}\pi^{-}$ or $D^{0}\to K^{+}K^{-}$, are described in terms of just two of the four amplitudes in Eq. (1) and their magnitudes can be parametrized as
$1-a_{D}^{f}=\left|\frac{\mathcal{A}_{\overline{D}}^{f}}{\mathcal{A}_{D}^{f}}\right|,$ (23)
where $a_{D}^{f}$ is the direct CP asymmetry, already introduced in Eq. (20). CP-conserving and CP-violating ratios of time-dependent decay rates are measured. The first type is obtained by calibrating the signal through the CF mode $K^{\mp}\pi^{\pm}$ as
$\frac{\Gamma(D^{0}\to f,t)+\Gamma(\overline{D^{0}}\to f,t)}{\Gamma(D^{0}\to K^{-}\pi^{+},t)+\Gamma(\overline{D^{0}}\to K^{+}\pi^{-},t)}\propto 1-\tau\tilde{y}_{\text{CP}},$ (24)
where we have neglected the quadratic terms in the charm mixing and CP-violating parameters that come with higher powers in $\tau$. Then, the slope $\tilde{y}_{\text{CP}}$ is extracted from Eq. (24) through a linear fit Aitala et al. (1999); Aaij et al. (2022); Link et al. (2000); Csorna et al. (2002); Zupanc et al. (2009); Lees et al. (2013); Ablikim et al. (2015); Starič et al. (2016); Aaij et al. (2019b); Nayak et al. (2020), and we have:
$2\tilde{y}_{\text{CP}}=y\cos\phi_{2}$ $\bigg{(}\left|\frac{q}{p}\right|+\left|\frac{p}{q}\right|\bigg{)}-\ x\sin\phi_{2}\bigg{(}\left|\frac{q}{p}\right|-\left|\frac{p}{q}\right|\bigg{)}$ (25)
-- $r_{D}^{K\pi}\bigg{[}\cos\phi_{2}(y\cos\delta_{D}^{K\pi}+x\sin\delta_{D}^{K\pi})\bigg{(}\left|\frac{q}{p}\right|+\left|\frac{p}{q}\right|\bigg{)}-\sin\phi_{2}(x\cos\delta_{D}^{K\pi}-y\sin\delta_{D}^{K\pi})\bigg{(}\left|\frac{q}{p}\right|-\left|\frac{p}{q}\right|\bigg{)}\bigg{]}.$
CP-violating ratios can be defined through the time-dependent asymmetries between the rates of $D^{0}$ and $\overline{D^{0}}$ decays to CP-even final states as follows:
$A^{\text{CP}}_{D}(f,t)=\frac{\Gamma(D^{0}\to f,t)-\Gamma(\overline{D^{0}}\to f,t)}{\Gamma(D^{0}\to f,t)+\Gamma(\overline{D^{0}}\to f,t)}=a_{D}^{f}+\tau\Delta Y^{f}.$ (26)
Experiments measure the slopes $\Delta Y^{f}$ of Eq. (26) for the modes $D^{0}\to K^{+}K^{-}$ and $D^{0}\to\pi^{+}\pi^{-}$aij et al. (2021a), through the following combinations
$\Delta Y^{KK}-$ $\ \Delta Y^{\pi\pi},$ (27)
$\Delta Y=\frac{1}{2}(\Delta Y^{KK}$ $+\Delta Y^{\pi\pi}).$
We use the relations in Eq. (27) to constrain the parameter $\Delta Y^{KK}$, which enters the observables coming from the WS/RS ratios of $D^{0}\to K^{\pm}\pi^{\mp}$ in Eq. (20). Moreover, since in the U-spin symmetric limit the final state-dependent contributions to $\Delta Y^{f}$ are equal and opposite in sign for $K^{+}K^{-}$ and $\pi^{+}\pi^{-}$ Kagan and Silvestrini (2021), averaging the two modes allows $\Delta Y$ to be expressed in the framework of approximate universality as
$\Delta Y=\frac{1}{2}\bigg{[}x\sin\phi_{2}\bigg{(}\left|\frac{q}{p}\right|+\left|\frac{p}{q}\right|\bigg{)}-y\cos\phi_{2}\bigg{(}\left|\frac{q}{p}\right|-\left|\frac{p}{q}\right|\bigg{)}\bigg{]},$ (28)
up to quadratic corrections in the charm mixing and CPV parameters.
Notice that we do not include the latest measurement of $\Delta Y^{f}$ from $D^{0}\to\pi^{+}\pi^{-}\pi^{0}$ decays by LHCb Aaij et al. (2024g) in the combination, as it is less precise than the observables in Eq. (27). Moreover, the application of approximate universality in this case is not as good as for $\Delta Y$ since there is no U-spin cancellation of final-state dependent contributions.
We consider in the combination also measurements of the time-integrated CP asymmetry for the mode $K^{+}K^{-}$ Aaij et al. (2023a) which, from Eq. (26), can be written as
$A_{D}^{\text{CP}}(f)=a_{D}^{f}+\left\langle\tau\right\rangle^{f}_{E}\Delta Y^{f},$ (29)
where we introduced the so-called average decay time $\left\langle\tau\right\rangle^{f}_{E}$, which depends on the final state $f$ and on the experimental environment $E$.
The first signal of direct CP violation in the charm system Aaij et al. (2019a) was observed in the difference between the asymmetries in Eq. (29) for the modes $K^{+}K^{-}$ and $\pi^{+}\pi^{-}$, which we fit as
$\Delta A^{\text{CP}}_{D}=a_{D}^{KK}-a_{D}^{\pi\pi}+\left\langle\tau\right\rangle^{KK}_{E^{\prime}}\Delta Y^{KK}-\left\langle\tau\right\rangle_{E^{\prime}}^{\pi\pi}\Delta Y^{\pi\pi}.$ (30)
III.3 Multi-body final states
The most precise observables for multi-body final states constraining together the charm mixing and CP-violating parameters come from the Dalitz plot analyses of $D^{0}\to K^{0}_{S}\pi^{+}\pi^{-}$ decays Aaij et al. (2019c, 2021b, 2023b). The reconstruction of the two-dimensional phase space and the decay time is performed, in a model-independent way, by partitioning them into bins, which we indicate as ±$\pm i$ and $j$, respectively, with $i$ and $-i$ connected by a CP transformation. For each of the bins, ratios analogous to WS/RS in Eq. (15) are measured:
$\frac{\Gamma_{-ij}(D^{0}\to K^{0}_{S}\pi^{+}\pi^{-})}{\Gamma_{ij}(D^{0}\to K^{0}_{S}\pi^{+}\pi^{-})},\qquad\qquad\frac{\Gamma_{ij}(\overline{D^{0}}\to K^{0}_{S}\pi^{+}\pi^{-})}{\Gamma_{-ij}(\overline{D^{0}}\to K^{0}_{S}\pi^{+}\pi^{-})}.$ (31)
Fitting the quantities in Eq. (31) with a quadratic expansion in $\tau$ of the decay rates, the following observables are determined:
$2x_{\text{CP}}=x\cos\phi_{2}\bigg{(}\left|\frac{q}{p}\right|+\left|\frac{p}{q}\right|\bigg{)}+y\sin\phi_{2}\bigg{(}\left|\frac{q}{p}\right|-\left|\frac{p}{q}\right|\bigg{)},\ \ \ \ 2y_{\text{CP}}=y\cos\phi_{2}\bigg{(}\left|\frac{q}{p}\right|+\left|\frac{p}{q}\right|\bigg{)}-x\sin\phi_{2}\bigg{(}\left|\frac{q}{p}\right|-\left|\frac{p}{q}\right|\bigg{)},$ (32)
[…]
Besides the Dalitz plot analysis, the phase space dependence of the $D$ mesons decay amplitudes to a multi-body final state $f$, such as $K^{0}_{S}K^{-}\pi^{+}$, can be parametrized through the following two quantities:
$(r_{D}^{f})^{2}=\ \frac{\int\text{d}\Phi_{D}\left|\mathcal{A}_{\overline{D}}^{f}\right|^{2}}{\int\text{d}\Phi_{D}\left|\mathcal{A}_{D}^{f}\right|^{2}}=$ $\frac{\int\text{d}\Phi_{D}\left|\mathcal{A}_{D}^{\overline{f}}\right|^{2}}{\int\text{d}\Phi_{D}\left|\mathcal{A}_{\overline{D}}^{\overline{f}}\right|^{2}},$ (33)
$\kappa_{D}^{f}e^{-i\delta_{D}^{f}}=\frac{\int\text{d}\Phi_{D}\mathcal{A}_{\overline{D}}^{f}\mathcal{A}_{D}^{f*}}{\sqrt{\int\text{d}\Phi_{D}\left|\mathcal{A}_{D}^{f}\right|^{2}}\sqrt{\int\text{d}\Phi_{D}\left|\mathcal{A}_{\overline{D}}^{f}\right|^{2}}}=$ $\frac{\int\text{d}\Phi_{D}\mathcal{A}_{D}^{\overline{f}}\mathcal{A}_{\overline{D}}^{\overline{f}*}}{\sqrt{\int\text{d}\Phi_{D}\left|\mathcal{A}_{\overline{D}}^{\overline{f}}\right|^{2}}\sqrt{\int\text{d}\Phi_{D}\left|\mathcal{A}_{D}^{\overline{f}}\right|^{2}}},$
where direct CP violation has been neglected. The integrals in Eq. (33) are performed over the $D$ phase space coordinates $\Phi_{D}$. The parameter in the upper row is a straightforward generalization for multi-body final states of the ratios introduced in Eqs. (14) and (23) for two-body decays. The interference integral in the bottom row is described simply through its magnitude $\kappa_{D}^{f}$, also known as the coherence factor, and its integrated strong phase $\delta_{D}^{f}$. For quasi-CP eigenstates, such as $\pi^{+}\pi^{-}\pi^{0}$, the imaginary part of the interference integral in Eq. (33) is vanishing. Then, instead of the coherence factor and the integrated strong phase, it is customary to use one single real parameter: the CP-even fraction $F_{+}^{f}$, defined as
$F^{f}_{+}=\frac{1}{2}\bigg{[}1+\kappa_{D}^{f}\cos\delta^{f}_{D}\bigg{]},$ (34)
which has already been introduced in Eq. (22). From its definition in Eq. (34), the CP-even fraction is non-negative and it is unity for two-body states.
Figure 1: Schematic representation of $B\to D$ cascade decays to two-body final states. The CKM angle $\gamma$ is extracted from the interference between the green and the red paths, whose amplitudes must be summed coherently. See the text for the definitions of the parameters and a generalization for multi-body final states.
The decay parameters for $K^{0}_{S}K^{\pm}\pi^{\mp}$ final states are measured from the ratio between the branching fractions of the DCS and CF decays Insler et al. (2012); Aaij et al. (2016b), given by
$\frac{\mathcal{B}(D^{0}\to K^{0}_{S}K^{+}\pi^{-})}{\mathcal{B}(D^{0}\to K^{0}_{S}K^{-}\pi^{+})}=\frac{(r_{D}^{K^{0}_{S}K\pi})^{2}-\kappa_{D}^{K^{0}_{S}K\pi}r_{D}^{K^{0}_{S}K\pi}(y\cos\delta_{D}^{K^{0}_{S}K\pi}-x\sin\delta_{D}^{K^{0}_{S}K\pi})}{1-\kappa_{D}^{K^{0}_{S}K\pi}r_{D}^{K^{0}_{S}K\pi}(y\cos\delta_{D}^{K^{0}_{S}K\pi}+x\sin\delta_{D}^{K^{0}_{S}K\pi})},$ (35)
where we have neglected CP violation and the non-linear terms in the mixing parameters.
IV Beauty observables
GLW Observables ADS Observables $\mathbf{B}\to\mathbf{D^{0}}$ decays $\mathbf{D^{0}}$ decays Ref.
$A_{B}^{\text{CP}}(XX\pi^{0},h)$, $A^{\text{fav,sup}}(K\pi\pi^{0},h)$, $B^{\pm}\to Dh^{\pm}$ $D\to K\pi\pi^{0}$, Aaij et al. (2015a)
$R^{\text{CP}}(XX\pi^{0},K\pi\pi^{0},K,\pi)$ $R^{\text{ADS}}(K\pi\pi^{0},h)$ $D\to XX\pi^{0}$
$A_{B}^{\text{CP}}(XX,h\pi\pi)$, $A^{\text{fav}}(K\pi,h\pi\pi)$, $B^{\pm}\to Dh^{\pm}\pi\pi$ $D\to K\pi$, Aaij et al. (2015b)
$R^{\text{CP}}(XX,K\pi,K\pi\pi,\pi\pi\pi)$ $R_{\pm}(K\pi,h\pi\pi)$ $D\to XX$
$A_{B}^{\text{CP}}(4\pi,h)$, $A^{\text{fav,sup}}(K3\pi,h)$, $B^{\pm}\to Dh^{\pm}$ $D\to K3\pi$, Aaij et al. (2016d)
$R^{\text{CP}}(4\pi,K3\pi,K,\pi)$ $R^{\text{ADS}}(K3\pi,h)$ $D\to 4\pi$
$A_{B}^{\text{CP}}(XX,K^{*\pm})$, $R(XX,K\pi,K^{*\pm})$ $A^{\text{fav,sup}}(K\pi,K^{*\pm})$, $R^{\text{ADS}}(K\pi,K^{*\pm})$ $B^{\pm}\to DK^{*\pm}$ $D\to K\pi$, $D\to XX$, Aaij et al. (2024c)
$A_{B}^{\text{CP}}(4\pi,K^{*\pm})$, $R(4\pi,K3\pi,K^{*\pm})$ $A^{\text{fav,sup}}(K3\pi,K^{*\pm})$, $R^{\text{ADS}}(K3\pi,K^{*\pm})$ $D\to K3\pi$, $D\to 4\pi$
$A_{B}^{\text{CP}}(XX,K^{*0})$, $R(XX,K\pi,K^{*0})$ $A^{\text{fav}}(K\pi,K^{*0})$, $R_{\pm}(K\pi,K^{*0})$ $B^{0}\to DK^{*0}$ $D\to K\pi$, $D\to XX$, Aaij et al. (2019d),
$A_{B}^{\text{CP}}(4\pi,K^{*0})$, $R(4\pi,K3\pi,K^{*0})$ $A^{\text{fav}}(K3\pi,K^{*0})$, $R_{\pm}(K3\pi,K^{*0})$ $B^{0}_{s}\to D\overline{K^{*0}}$ $D\to K3\pi$, $D\to 4\pi$ Aaij et al. (2024b)
$A_{B}^{\text{CP}}(XX,h)$, $R^{\text{CP}}(XX,K\pi,K,\pi)$ $A^{\text{fav}}(K\pi,K)$, $R_{\pm}(K\pi,h)$ $B^{\pm}\to Dh^{\pm}$ $D\to K\pi$, $D\to XX$ Aaij et al. (2021c)
$A_{B}^{\text{CP}}(XX,h(\gamma,\pi^{0}))$, $A^{\text{fav}}(K\pi,K(\gamma,\pi^{0}))$, $B^{\pm}\to[D\pi^{0}]_{D^{*}}h^{\pm}$ $D\to K\pi$, Aaij et al. (2021c)
$R^{\text{CP}}(XX,K\pi,K(\gamma,\pi^{0}),\pi(\gamma,\pi^{0}))$ $R_{\pm}(K\pi,h(\gamma,\pi^{0}))$ $B^{\pm}\to[D\gamma]_{D^{*}}h^{\pm}$ $D\to XX$
$A^{\text{fav,sup}}(K^{0}_{S}K\pi,h)$,
—— $R^{\text{fav,sup}}(K^{0}_{S}K\pi,K,\pi)$, $B^{\pm}\to Dh^{\pm}$ $D\to K^{0}_{S}K\pi$ Aaij et al. (2020),
$R^{\text{ADS}}(K^{0}_{S}K\pi,\pi)$ Adachi et al. (2023)
$A_{B}^{\text{CP}}(XX\pi\pi,h)$, $R^{\text{CP}}(XX\pi\pi,K3\pi,K,\pi)$ —— $B^{\pm}\to Dh^{\pm}$ $D\to XX\pi\pi$, $D\to K3\pi$ Aaij et al. (2023c)
$R^{\text{ADS}}(K\pi,h)$, $A^{\text{sup}}(K\pi,h)$ $B^{\pm}\to Dh^{\pm}$ $D\to K\pi$ Horii et al. (2011)
$R^{\text{ADS}}(K\pi\pi^{0},h)$, $A^{\text{sup}}(K\pi\pi^{0},h)$ $B^{\pm}\to Dh^{\pm}$ $D\to K\pi\pi^{0}$ Nayak et al. (2013)
$A_{B}^{\text{CP}}(K^{+}K^{-},K)$, $R^{\text{CP}}(K^{+}K^{-},K\pi,K,\pi)$ $B^{\pm}\to Dh^{\pm}$ $D\to K^{+}K^{-}$, $D\to K\pi$, Adachi et al. (2024a)
$A_{B}^{\text{CP}}(K^{0}_{S}\pi^{0},K)$ , $R^{\text{CP}}(K^{0}_{S}\pi^{0},K\pi,K,\pi)$ —— $D\to K^{0}_{S}\pi^{0}$
$A_{B}^{\text{CP}}(f^{\pm},K)$, $R^{\text{CP}}(f^{\pm},K\pi,K,\pi)$ $B^{\pm}\to Dh^{\pm}$ $D\to f^{\pm}$, Abe et al. (2006)
$A_{B}^{\text{CP}}(f^{\pm},K(\pi^{0}))$, —— $B^{\pm}\to[D\pi^{0}]_{D^{*}}h^{\pm}$ $D\to K\pi$
$R^{\text{CP}}(f^{\pm},K\pi,K(\pi^{0}),\pi(\pi^{0}))$
Table 2: GLW and ADS observables coming from $B$ mesons cascade decays. The $XX$ notation indicate both $K^{+}K^{-}$ and $\pi^{+}\pi^{-}$ states, while we use $K\pi$ for $K^{\mp}\pi^{\pm}$. $h$ stands for both $K$ and $\pi$ mesons. We refer to $\pi^{+}\pi^{-}\pi^{+}\pi^{-}$ and $K^{\mp}\pi^{\pm}\pi^{+}\pi^{-}$ final states as $4\pi$ and $K3\pi$, respectively. We use $f^{\pm}$ as a shorthand notation for CP-even and CP-odd eigenstates, respectively. The observables are described in details in Apps. A and B.
The precision on the charm mixing and decay parameters can be improved by including in the combination time-integrated measurements of the so-called $B$ meson cascade decays to $D^{0}-\overline{D^{0}}$ mixed states Rama (2014); Aaij et al. (2021f); LHC (2024). We show how the CKM angle $\gamma$ appears in the equations describing the most used and precise observables available to date.
We discuss the extraction of $\gamma$ from the CP-violating phase of the interference between neutral $B$ decays to charmed mesons with and without mixing that is used to test the consistency of the combination.
IV.1 Cascade decays
Beauty cascade decays Abe et al. (2006); Horii et al. (2011); Nayak et al. (2013); Aaij et al. (2015a, b, 2016d); LHC (2016); Aaij et al. (2016e, 2019d, 2021c, 2020); Adachi et al. (2023, 2024a); Aaij et al. (2024b, c) are processes where a $B$ meson goes to an hadron state $h$ and a $D^{0}$ or $\overline{D^{0}}$ meson, which subsequently oscillates and decays to one of the final states $f$ introduced in Sec. III. In the following, we indicate with $B$ a negatively charged $B^{-}$ or a neutral meson $\overline{B^{0}_{(s)}}$.
Since the final states $f$ are accessible to both the neutral $D$ mesons, the total amplitude for the cascade decays is given by the coherent sum of the two paths in Fig. 1. Then, the CKM angle $\gamma$ can be measured from the interference between the favoured $b\to c$ and suppressed $b\to u$ quark transitions and the corresponding decay amplitudes can be arranged as
$(r_{B}^{Dh})^{2}=\frac{\int\text{d}\Phi_{B}\left|\mathcal{A}_{B}^{\overline{D}h}\right|^{2}}{\int\text{d}\Phi_{B}\left|\mathcal{A}_{B}^{Dh}\right|^{2}},\qquad\qquad\kappa_{B}^{Dh}e^{i(\delta_{B}^{Dh}-\gamma)}=\frac{\int\text{d}\Phi_{B}\mathcal{A}_{B}^{\overline{D}h}(\mathcal{A}_{B}^{Dh})^{*}}{\sqrt{\int\text{d}\Phi_{B}\left|\mathcal{A}_{B}^{Dh}\right|^{2}}\sqrt{\int\text{d}\Phi_{B}\left|\mathcal{A}_{B}^{\overline{D}h}\right|^{2}}},$ (36)
with the integrals performed over the $B$ meson phase space $\Phi_{B}$. For two-body final states, the expressions in Eq. (36) simplify since there is no phase space dependence and the coherence factor is unity.
Then, the time-integrated rate of the cascade decay $B\to[f]_{D}h$ is found to be
$\Gamma(B\to[f]_{D}h)\propto 1+(r^{f}_{D}r_{B}^{Dh})^{2}+2\kappa_{D}^{f}\kappa_{B}^{Dh}r_{D}^{f}r_{B}^{Dh}\cos(\delta_{B}^{Dh}-\delta_{D}^{f}-\gamma)+\Gamma_{\mathrm{mix}}(x,y,\gamma,f,h),$ (37)
where $\Gamma_{\mathrm{mix}}(x,y,\gamma,f,h)$ takes into account $D^{0}-\overline{D^{0}}$ mixing and it is given by
$\Gamma_{\mathrm{mix}}(x,y,\gamma,f,h)=$ $\ -\alpha y\bigg{[}\kappa_{D}^{f}r_{D}^{f}\cos\delta_{D}^{f}(1+(r_{B}^{Dh})^{2})+\kappa_{B}^{Dh}r_{B}^{Dh}\cos(\delta_{B}^{Dh}-\gamma)(1+(r_{D}^{f})^{2})\bigg{]}$ (38)
$\ +\alpha x\bigg{[}\kappa_{D}^{f}r_{D}^{f}\sin\delta_{D}^{f}((r_{B}^{Dh})^{2}-1)+\kappa_{B}^{Dh}r_{B}^{Dh}\sin(\delta_{B}^{Dh}-\gamma)(1-(r_{D}^{f})^{2})\bigg{]},$
up to quadratic corrections in the charm mixing and CP-violating parameters and neglecting direct CP violation.
GGSZ Observables $\mathbf{B}\to\mathbf{D^{0}}$ decays $\mathbf{D^{0}}$ decays Ref.
$x_{\pm}^{DK^{*0}}$, $y_{\pm}^{DK^{*0}}$ $B^{0}\to DK^{*0}$ $D\to K^{0}_{S}XX$ Aaij et al. (2016f, 2024e)
$x^{DK}_{\pm}$, $x^{D\pi}_{\xi}$ $B^{\pm}\to DK^{\pm}$ $D\to K^{0}_{S}XX$ Resmi et al. (2019); Aaij et al. (2021d); Abudinén et al. (2022)
$y_{\pm}^{DK}$, $y^{D\pi}_{\xi}$ $B^{\pm}\to D\pi^{\pm}$
$x^{DK}_{\pm}$, $x^{D\pi}_{\xi}$ $B^{\pm}\to DK^{\pm}$ $D\to XX\pi\pi$ Aaij et al. (2023c)
$y_{\pm}^{DK}$, $y^{D\pi}_{\xi}$ $B^{\pm}\to D\pi^{\pm}$
$x^{[D\pi^{0}]_{D^{*}}K}_{\pm}$, $x^{[D\gamma]_{D^{*}}K}_{\pm}$ $B^{\pm}\to[D\pi^{0}]_{D^{*}}K^{\pm}$ $D\to K^{0}_{S}\pi\pi$ Poluektov et al. (2010)
$y_{\pm}^{[D\pi^{0}]_{D^{*}}K}$, $y^{[D\gamma]_{D^{*}}K}_{\pm}$ $B^{\pm}\to[D\gamma]_{D^{*}}K^{\pm}$
$x^{D^{*}K}_{\pm}$, $x^{D^{*}\pi}_{\xi}$ $B^{\pm}\to D^{*}K^{\pm}$ $D\to K^{0}_{S}XX$ Aaij et al. (2024d, 2023d)
$y_{\pm}^{D^{*}K}$, $y^{D^{*}\pi}_{\xi}$ $B^{\pm}\to D^{*}\pi^{\pm}$
$x_{\pm}^{DK^{*\pm}}$, $y_{\pm}^{DK^{*\pm}}$ $B^{\pm}\to DK^{*\pm}$ $D\to K^{0}_{S}XX$ Aaij et al. (2024c)
Table 3: GGSZ observables used in the combination. Here, $XX$ stands for both the $K^{+}K^{-}$ and $\pi^{+}\pi^{-}$ states. The dependence of the observables on the fit parameters are reported in Eqs. (39) and (41).
The $\alpha$ coefficient in Eq. (38) describes the impact on $\Gamma_{\mathrm{mix}}(x,y,\gamma,f,h)$ of the non-trivial dependence of the signal selection efficiency on time and it is one when the latter is constant Rama (2014).
Since at this order in the charm mixing and CP-violating parameters the decay rate of the CP conjugated process is obtained simply by replacing $\gamma$ with $-\gamma$ in Eq. (37), the difference between $\Gamma(B\to[f]_{D}h)$ and $\Gamma(\overline{B}\to[\overline{f}]_{D}\overline{h})$ gives access to $\sin\gamma$. This is commonly referred to as the time-integrated CP asymmetry, which is an example of the so-called Gronau-London-Wyler (GLW) Gronau and Wyler (1991); Gronau and London (1991) and Atwood-Dunietz-Soni (ADS) Atwood et al. (1997, 2001) observables that we use in our combination. A complete list of the GLW and ADS modes considered in this work is reported in Tab. 2, while their parametrizations are derived in Apps. A and B. We have taken the coherence factors for the modes $B\to DK^{*}$ from LHC (2016) and Aaij et al. (2016e) for charged and neutral $B$ mesons, respectively.
For cascade decays with multi-body final states of the charm, as $K^{0}_{S}K^{+}K^{-}$$(\pi^{+}\pi^{-})$ or $\pi^{+}\pi^{-}K^{+}K^{-}$$(\pi^{+}\pi^{-})$, analyses of the variation of the decay rates across the phase space are available Aaij et al. (2016f); Resmi et al. (2019); Aaij et al. (2021d); Abudinén et al. (2022); Aaij et al. (2023c, 2024d, d, 2024e); Poluektov et al. (2010); Aaij et al. (2024c). The Dalitz plots of these events are fitted using a model for the $D$ meson decay amplitudes or by binning the phase space and solving a system of linear equations in a model-independent way. The first method is more precise statistically with respect to the latter but introduces an additional systematic uncertainty due to the model. In both cases, the so-called Giri-Grossmann-Soffer-Zupan (GGSZ) Giri et al. (2003) observables are obtained, depending on $\gamma$ as
$x_{\pm}^{Dh}=r_{B}^{Dh}\cos(\delta_{B}^{Dh}\pm\gamma),\qquad\qquad y_{\pm}^{Dh}=r_{B}^{Dh}\sin(\delta_{B}^{Dh}\pm\gamma).$ (39)
Measuring the quantities in Eq. (39) for beauty decays with a pion in the final state is particularly hard due to the small value of $r_{B}^{D\pi}\sim\mathcal{O}(10^{-3})$. In this case, the GGSZ observables for $B\to[f]_{D}\pi$ are written as rotated with respect to the ones entering $B\to[f]_{D}K$ through the transformation
$\begin{pmatrix}x_{\pm}^{D\pi}\\ y_{\pm}^{D\pi}\end{pmatrix}=\begin{pmatrix}x_{\xi}^{D\pi}&-y_{\xi}^{D\pi}\\ y_{\xi}^{D\pi}&x_{\xi}^{D\pi}\end{pmatrix}\begin{pmatrix}x_{\pm}^{DK}\\ y_{\pm}^{DK}\end{pmatrix},$ (40)
where $x_{\xi}^{D\pi}$ and $y_{\xi}^{D\pi}$ are given by
$x_{\xi}^{D\pi}=\frac{r_{B}^{D\pi}}{r_{B}^{DK}}\cos(\delta_{B}^{D\pi}-\delta_{B}^{DK}),\qquad\qquad y_{\xi}^{D\pi}=\frac{r_{B}^{D\pi}}{r_{B}^{DK}}\sin(\delta_{B}^{D\pi}-\delta_{B}^{DK}).$ (41)
Then, the coefficients in Eq. (41) are extracted simultaneously with $x_{\pm}^{DK}$ and $y_{\pm}^{DK}$. The GGSZ observables used in our combination are reported in Tab. 3.
It is worth noting that the expressions in Eqs. (37), (38), (39) and (41) are still valid for cascade decays with the $D^{0}$ or $\overline{D^{0}}$ meson produced together with a $\pi^{0}$ meson or a photon $\gamma$ by the decay of a $D^{*}$ intermediate state, simply replacing the strong phase $\delta_{D}^{f}$ with $\delta_{D^{*}}^{f}$ or $\delta_{D^{*}}^{f}+\pi$, respectively.
IV.2 Neutral $B$ decays to charmed mesons
Figure 2: Schematic representation of the extraction of the CKM angle $\gamma$ from the CP-violating phase of the interference between neutral $B$ decays to charmed mesons with and without mixing. Here, $h^{+}$ stands for $\pi^{+}$ or $K^{+}(\pi^{+}\pi^{-})$ for $B^{0}$ and $B^{0}_{s}$ mesons, respectively.
Observables $\mathbf{B^{0}_{(s)}}\to\mathbf{D_{(s)}^{\mp}}\mathbf{h^{\pm}}$ decays Ref.
$C_{D_{s}^{\mp}K^{\pm}}$,           $G_{D_{s}^{\mp}K^{\pm}}$,           $S_{D_{s}^{\mp}K^{\pm}}$ $B^{0}_{s}\to D_{s}K$ Aaij et al. (2024f)
$S_{D^{\mp}\pi^{\pm}}$ $B^{0}\to D\pi$ Aaij et al. (2018)
$C_{D_{s}^{\mp}K^{\pm}\pi\pi}$,          $G_{D_{s}^{\mp}K^{\pm}\pi\pi}$,          $S_{D_{s}^{\mp}K^{\pm}\pi\pi}$ $B^{0}_{s}\to D_{s}K\pi\pi$ Aaij et al. (2021e)
Table 4: Observables coming from time-dependent analyses of neutral $B$ mesons decays to charmed mesons. The observables are defined in Eq. (44).
The CKM angle $\gamma$ can be obtained from the CP-violating phase of the interference between neutral $B$ decays to charmed mesons $f$ (e.g. $f=D^{-}\pi^{+}$ or $f=D^{-}_{s}K^{+}$$(\pi^{+}\pi^{-})$) with and without mixing, as shown schematically in Fig. 2. The decay amplitudes for these processes can be organized into ratios of magnitudes, coherence factors and strong phases through the following integrals over the phase space $\Phi_{B^{0}_{(s)}}$:
$(r_{B^{0}_{(s)}}^{f})^{2}=\frac{\int\text{d}\Phi_{B^{0}_{(s)}}\left|\mathcal{A}_{\overline{B^{0}_{(s)}}}^{f}\right|^{2}}{\int\text{d}\Phi_{B^{0}_{(s)}}\left|\mathcal{A}_{B^{0}_{(s)}}^{f}\right|^{2}},\qquad\qquad\kappa_{B^{0}_{(s)}}^{f}e^{i(\delta_{B^{0}_{(s)}}^{f}+\varphi_{(s)})}=\frac{\int\text{d}\Phi_{B^{0}_{(s)}}\mathcal{A}_{\overline{B^{0}_{(s)}}}^{f}\ \mathcal{A}_{B^{0}_{(s)}}^{f*}}{\sqrt{\int\text{d}\Phi_{B^{0}_{(s)}}\left|\mathcal{A}_{\overline{B^{0}_{(s)}}}^{f}\right|^{2}}\sqrt{\int\text{d}\Phi_{B^{0}_{(s)}}\left|\mathcal{A}_{B^{0}_{(s)}}^{f}\right|^{2}}},$ (42)
where $\varphi_{(s)}$ are the weak phases entering the decay, which are given in the SM by a combination of the relevant CKM matrix elements. Then, the time-dependent rates of neutral $B^{0}_{(s)}$ meson decays can be written as Dunietz et al. (2001); Artuso et al. (2016)
$\Gamma(B^{0}_{(s)}\to f,t)\propto\cosh(t\Delta\Gamma_{(s)}/2)-G_{f}\sinh(t\Delta\Gamma_{(s)}/2)+C_{f}\cos(t\Delta m_{(s)})-S_{f}\sin(t\Delta m_{(s)}),$ (43)
where $\Delta m_{(s)}$ and $\Delta\Gamma_{(s)}$ are the differences in masses and lifetimes between the heavier and lighter eigenstates of the $B^{0}_{(s)}$ mixing Hamiltonian. The observables $C_{f}$, $G_{f}$ and $S_{f}$ are fitted from the data Aaij et al. (2018, 2021e, 2024f) and are given by
$C_{f}=\frac{1-(r_{B^{0}_{(s)}}^{f})^{2}}{1+(r_{B^{0}_{(s)}}^{f})^{2}},\ \ \ \ \ G_{f}=\frac{-2\kappa_{B^{0}_{(s)}}^{f}r_{B^{0}_{(s)}}^{f}}{1+(r_{B^{0}_{(s)}}^{f})^{2}}\cos(\delta_{B^{0}_{(s)}}^{f}+(\phi_{(s)}-\gamma)),\ \ \ \ \ S_{f}=\frac{2\kappa_{B^{0}_{(s)}}^{f}r_{B^{0}_{(s)}}^{f}}{1+(r_{B^{0}_{(s)}}^{f})^{2}}\sin(\delta_{B^{0}_{(s)}}^{f}+(\phi_{(s)}-\gamma)).$ (44)
Here, we have neglected CP violation in pure $B^{0}_{(s)}-\overline{B^{0}_{(s)}}$ mixing, as well as direct CP violation.
The phases $\phi_{(s)}$ are obtained in the SM from the CKM angles $\beta=\arg(-V_{td}^{*}V_{cd}V_{cb}^{*}V_{tb})$ and $\beta_{s}=\arg(-V_{cs}^{*}V_{ts}V_{tb}^{*}V_{cb})$ as $\phi=-2\beta+\pi$ and $\phi_{s}=2\beta_{s}$, up to corrections to the fourth power of the sine of the Cabibbo angle.
Figure 3: P.d.f.s of the charm mixing and CP-violating parameters and the CKM angle $\gamma$ obtained using all the observables. Darker (lighter) regions correspond to $68.3\%$ ($95.4\%$) probability.
Beyond the SM, the relations between $\phi_{(s)}$ and $\beta_{(s)}$ could be altered, but one can still use the experimental values of $\phi_{(s)}$ to take into account $B^{0}_{(s)}$ mixing effects in the extraction of $\gamma$. The measurements used in the combination are reported in Tab. 4, while inputs for $\beta$ and $\beta_{s}$ are taken from the world averages in Amhis et al. (2023) and Zyla et al. (2020), respectively. Similar observables to the ones in Eq. (44) can be obtained for the CP conjugate final states by changing the signs of $C_{f}$, $S_{f}$ and of the CP-violating phases between mixing and decay $(\phi_{(s)}-\gamma)$.
V Results
Par. Value Unc. 95.4% Prob. 99.7% Prob.
$x_{12}\simeq x[\%]$ $0.402$$\pm 0.034$ $[0.335,0.470]$ $[0.301,0.503]$
$y_{12}\simeq y[\%]$ $0.629$$\pm 0.021$ $[0.587,0.672]$ $[0.566,0.693]$
$\phi_{2}^{M}[\degree]$ $2.4$$\pm 1.6$ $[-0.7,5.7]$ $[-2.3,7.5]$
ϕ2Γ[°]superscriptsubscriptitalic-$\phi_{2}^{\Gamma}[\degree]$ $2.6$$\pm 1.2$ $[0.2,5.0]$ $[-1.0,6.2]$
ϕ12[°]subscriptitalic-$\phi_{12}[\degree]$ $-0.2$$\pm 2.0$ $[-4.1,4.0]$ $[-6.1,6.1]$
ϕ2[°]subscriptitalic-$\phi_{2}[\degree]$ $-2.54$$\pm 0.94$ $[-4.44,-0.66]$ $[-5.41,0.28]$
$|q/p|-1[\%]$ $-0.1$$\pm 1.6$ $[-3.3,3.1]$ $[-4.8,4.7]$
$a_{D}^{KK}[\%]$ $0.085$$\pm 0.063$ $[-0.041,0.211]$ $[-0.105,0.273]$
$a_{D}^{\pi\pi}[\%]$ $0.301$$\pm 0.070$ $[0.161,0.443]$ $[0.090,0.513]$
$\gamma[\degree]$ $68.4$$\pm 2.6$ $[63.2,73.5]$ $[60.3,75.8]$
Table 5: Results of the combination when using all the observables. The half-width of the smallest $68.3\%$ probability interval (Unc.) and its central value (Value) are reported for each quantity of interest. The smallest intervals containing the $95.4\%$ and $99.7\%$ probabilities are reported in the last two columns on the right.
Figure 4: Two-dimensional contours of the charm mixing and CP-violating parameters obtained using all the observables. Darker (lighter) regions correspond to $68.3\%$ ($95.4\%$) probability.
We combine the observables in Tabs. 1, 2, 3 and 4 in a Bayesian framework, employing as likelihood the product of Gaussian distributions for each set of correlated measurements. The fit parameters are assumed to follow uniform distributions defined over the ranges in Tabs. 7 and 8. It is worth noting that the beauty observables described in Sec. IV and Apps. A and B are symmetric under the simultaneous transformation $\gamma\to\gamma\pm\pi$ and $\delta_{B}^{Dh}\to\delta_{B}^{Dh}\pm\pi$, for each of the $B$ modes considered. Thus, we have limited our study to just one of these solutions by defining the prior of $\gamma$ over a sub-set contained in $(0,\pi)$. Then, other solutions are found by adding ±$\pm\pi$ to our results. The posterior pdfs of the parameters are determined through a Markov Chain Monte Carlo (MCMC) algorithm [^1] The fit code is available at https://github.com/silvest/GammaDDbar Di Palma and Silvestrini (2024). implemented in the Bayesian Analysis Toolkit (BAT) software package Caldwell et al. (2009). We perform four combinations using different subsets of the beauty observables, divided by the species of the $B$ mesons: all the modes together, only charged $B$, only neutral $B$ and $B_{s}$ modes.
We report in Tab. 5 the charm mixing and CP-violating parameters and the CKM angle $\gamma$ obtained using all the inputs, while the corresponding pdfs are shown in Fig. 3. Two-dimensional contours are depicted in Fig. 4. The rest of the fit results are reported in Tabs. 10 and 11 and Fig. 6. Correlation between $D$ meson mixing parameters and the CKM angle $\gamma$ is below $1\%$, as reported in Tab. 9.
The charm parameters are compatible with the estimates found by the 2023 analysis by HFLAV [^2] See the “No subleading ampl. for CF/DCS decays” fit at https://hflav-eos.web.cern.ch/hflav-eos/charm/CKM23/results_mix_cpv.html. Amhis et al. (2023).
$\mathbf{B}$ meson types Value Unc. $\mathbf{95.4}\%$ Prob. $\mathbf{99.7}\%$ Prob.
$\mathbf{B^{\pm}:}\ \gamma_{B^{\pm}}[\degree]$ $67.4$$\pm 3.1$ $[61.0,73.3]$ $[57.6,76.0]$
$\mathbf{B^{0}_{s}:}\ \gamma_{B^{0}_{s}}[\degree]$ $76.8$$\pm 9.6$ $[57.9,95.9]$ $[49.5,105.1]$
$\mathbf{B^{0}:}\ \gamma_{B^{0}}[\degree]$ $69.5$$\pm 5.7$ $[56.6,80.4]$ $[47.1,85.3]$
Table 6: Probability intervals for $\gamma$ when splitting the inputs of the combination: only charged $B$ measurements ($\gamma_{B^{\pm}}$), only neutral $B$ ($\gamma_{B^{0}}$) and $B_{s}$ modes ($\gamma_{B^{0}_{s}}$). The central value (Value) and the half-width of the smallest $68.3\%$ probability interval (Unc.) are reported. The last two columns on the right provide the smallest intervals containing the $95.4\%$ and $99.7\%$ probabilities.
Figure 5: Pdfs of the three estimates of $\gamma$ obtained combining measurements involving only charged $B$ ($\gamma_{B^{\pm}}$), only $B^{0}$ ($\gamma_{B^{0}}$) or $B^{0}_{s}$ ($\gamma_{B^{0}_{s}}$) mesons.
Considering beauty observables in the combination allows for an improvement at the percent level in the precision of the mixing parameters $x$ and $y$ with respect to the HFLAV results. This can be understood by looking at the $D$ mixing contribution to $B$ decay rates $\Gamma_{\text{mix}}(x,y,\gamma,f,h)$ in Eq. (38), or at the expressions in Apps. A and B. They show that the contributions to beauty cascade decays arising due to $D^{0}-\overline{D^{0}}$ mixing range from $\mathcal{O}(10^{-3})$ to $\mathcal{O}(10^{-1})$, depending on the specific observable and decay channel. Thus, currently only the most precise beauty measurements, which have relative uncertainties comparable to the charm mixing contributions, can effectively constrain $x$ and $y$. However, the impact of the combined analysis of charm and beauty observables on the determination of the $D$ mixing parameters will increase significantly as $B$ cascade decays measurements improve in the future.
The CKM angle $\gamma$ falls within the confidence intervals found by the latest LHCb combination LHC (2024), as the other beauty and charm parameters. The $\gamma$ estimates extracted using measurements of beauty decays with different types of initial $B$ mesons separately are reported in Tab. 6 and their pdfs are shown in Fig. 5. Our results are all compatible with each other, substantially reducing the moderate tension of $2.2\sigma$ between $\gamma_{B^{\pm}}$ and $\gamma_{B^{0}}$, observed by the LHCb combination and in our previous work Di Palma (2024), where we found $\gamma_{B^{0}}=(81\pm 11)\degree$. This change is driven by the most precise measurement of $B^{0}\to DK^{*0}$ cascade decay recently published by LHCb Aaij et al. (2024b), which already in their work provided a $\gamma$ estimate compatible with our result.
VI Conclusions
We presented a combination of charm and beauty observables to determine the CKM angle $\gamma$ and the neutral $D$ mixing and CP-violating parameters simultaneously in the framework of the SM and employing approximate universality for $D^{0}-\overline{D^{0}}$ mixing.
The charm parameters are given by
$x_{12}\simeq x=(0.402\pm 0.034)\%,\qquad y_{12}\simeq y=(0.629\pm 0.021)\%,\qquad\qquad\phi_{2}^{M}=(2.4\pm 1.6)\degree,\qquad\phi_{2}^{\Gamma}=(2.6\pm 1.2)\degree,$
$\phi_{2}=(-2.54\pm 0.94)\degree,\qquad\qquad\left|q/p\right|-1=(-0.1\pm 1.6)\%,$
while the CKM angle $\gamma$ is found to be
$\gamma=(68.4\pm 2.6)\degree.$
The uncertainties on the two CP-violating phases $\phi_{2}^{M,\Gamma}$ are still one order of magnitude larger than what is needed to test the predictions $\phi_{2}^{M}\sim\phi_{2}^{\Gamma}\sim 0.1\degree$, obtained from U-spin arguments in Eq. (10), but a factor of four larger than the upper bound of $\left|\phi_{2}^{\Gamma}\right|<0.3\degree$ in Eq. (11). An interesting byproduct of our fit is the value of the direct CP asymmetries in $D^{0}\to K^{+}K^{-}$ and $D^{0}\to\pi^{+}\pi^{-}$ decays within the framework of approximate universality:
$a_{D}^{KK}=(8.5\pm 6.3)\cdot 10^{-4}\,,\qquad a_{D}^{\pi\pi}=(30.1\pm 7.0)\cdot 10^{-4}\,,\qquad\rho=67\%\,,$ (45)
providing evidence for direct CP violation in $D^{0}\to\pi^{+}\pi^{-}$ at more than $4\sigma$, together with a sizable deviation from the U-spin expectation of $a_{D}^{KK}\sim-a_{D}^{\pi\pi}$.
To check the consistency of our combination of measurements sensitive to the angle $\gamma$, we extracted the value of $\gamma$ separately from charged $B$, $B^{0}$ and $B^{0}_{s}$ meson decays. We found three compatible estimates, given by
$\gamma_{B^{\pm}}=(67.4\pm 3.1)\degree,\qquad\qquad\gamma_{B^{0}}=(69.5\pm 5.7)\degree,\qquad\qquad\gamma_{B^{0}_{s}}=(76.8\pm 9.6)\degree.$
Acknowledgements.
We warmly thank F. Blanc, A. Gilman, M. Morello and S. Vecchi for correspondence on LHCb results. This work was supported in part by the European Union - Next Generation EU under italian MUR grant PRIN-2022-RXEZCJ and by the HSE basic research fund.
Appendix A GLW Observables
The GLW observables consist of CP-conserving and CP-violating ratios of time-integrated rates of $B$ cascade decays when the final states $f$ are (quasi-)CP eigenstates.
Here, we report the definitions of the observables in the leftmost column of Tab. 2 and their expressions up to quadratic corrections in charm mixing and CP-violating parameters, neglecting direct CP violation.
The CP asymmetry $A_{B}^{\text{CP}}(f,h)$ is defined as
$A_{B}^{\text{CP}}(f,h)=\frac{\Gamma(B\to[f]_{D}h)-\Gamma(\overline{B}\to[f]_{D}\overline{h})}{\Gamma(B\to[f]_{D}h)+\Gamma(\overline{B}\to[f]_{D}\overline{h})}=\frac{2r_{B}^{Dh}\kappa_{B}^{Dh}\sin\gamma\sin\delta_{B}^{Dh}}{(1+(r_{B}^{Dh})^{2})\frac{1-(2F_{+}^{f}-1)\alpha y}{2F_{+}^{f}-1-\alpha y}+2r_{B}^{Dh}\kappa_{B}^{Dh}\cos\gamma\cos\delta_{B}^{Dh}}.$ (46)
The ratio of rates for decays in which the $D$ meson is reconstructed as a CP eigenstate and decays with CF final states $f^{*}$, as $K^{-}\pi^{+}(\pi^{0})$ or $K^{-}\pi^{+}\pi^{+}\pi^{-}$, is given by
$R(f,f^{*},h)=\frac{\mathcal{B}(D^{0}\to f^{*})}{\mathcal{B}(D^{0}\to f)}\frac{\Gamma(B\to[f]_{D}h)+\Gamma(\overline{B}\to[f]_{D}\overline{h})}{\Gamma(B\to[f^{*}]_{D}h)+\Gamma(\overline{B}\to[\overline{f^{*}}]_{D}\overline{h})}.$ (47)
In terms of charm parameters, Eq. (47) can be written as
$R(f,f^{*},h)=\frac{(1+(r_{B}^{Dh})^{2})(1-(2F_{+}^{f}-1)\alpha y)+2r_{B}^{Dh}\kappa_{B}^{Dh}\cos\gamma\cos\delta_{B}^{Dh}((2F_{+}^{f}-1)-\alpha y)}{1+(r_{D}^{f^{*}}r_{B}^{Dh})^{2}+2r_{B}^{Dh}\kappa_{B}^{Dh}r_{D}^{f^{*}}\kappa_{D}^{f^{*}}\cos\gamma\cos(\delta_{B}^{Dh}-\delta_{D}^{f^{*}})+\Gamma_{\mathrm{mix}}^{+}(x,y,\gamma,f^{*},h)},$ (48)
with $\Gamma^{+}_{\text{mix}}(x,y,\gamma,f^{*},h)$ arising due to the mixing terms in Eq. (38), given by
$\Gamma_{\mathrm{mix}}^{+}(x,y,\gamma,f^{*},h)=$ $\ -\alpha y\bigg{[}r_{D}^{f^{*}}\kappa_{D}^{f^{*}}(1+(r_{B}^{Dh})^{2})\cos\delta_{D}^{f^{*}}+r_{B}^{Dh}\kappa_{B}^{Dh}(1+(r_{D}^{f^{*}})^{2})\cos\gamma\cos\delta_{B}^{Dh}\bigg{]}$ (49)
$\ +\alpha x\bigg{[}r_{B}^{Dh}\kappa_{B}^{Dh}(1-(r_{D}^{f^{*}})^{2})\cos\gamma\sin\delta_{B}^{Dh}-r_{D}^{f^{*}}\kappa_{D}^{f^{*}}(1-(r_{B}^{Dh})^{2})\sin\delta_{D}^{f^{*}}\bigg{]}.$
Sometimes, it is useful to avoid the dependence on the charm branching fractions in Eq. (47) by measuring directly the so-called double ratio, which is related to $R(f,f^{*},h)$ for two different meson states $h$ and $h^{\prime}$ as
$R^{\text{CP}}(f,f^{*},h,h^{\prime})=\frac{R(f,f^{*},h)}{R(f,f^{*},h^{\prime})}.$ (50)
Appendix B ADS Observables
In this section, we give an overview of the ADS observables considered in the combination in Tab. 2, expanding them in terms of the charm parameters and neglecting the non-linear terms, as well as direct CP violation. The ADS observables are defined through time-integrated rates of $B$ cascade decays, involving CF/DCS decays of the $D$ mesons. As in Sec. III.1, we refer to the CF mode of the $D^{0}$ meson as $f$.
The simplest observables that can be measured are the CP conjugate suppressed over favoured ratios
$R^{-}(f,h)=\frac{\Gamma(B\to[\overline{f}]_{D}h)}{\Gamma(B\to[f]_{D}h)},\qquad\qquad R^{+}(f,h)=\frac{\Gamma(\overline{B}\to[f]_{D}\overline{h})}{\Gamma(\overline{B}\to[\overline{f}]_{D}\overline{h})}.$ (51)
The rates for the favoured processes are reported in Eq. (37), while for the suppressed modes, we have
$\Gamma(B\to[\overline{f}]_{D}h)\propto(r_{D}^{f})^{2}+(r_{B}^{Dh})^{2}+2r_{B}^{Dh}r_{D}^{f}\kappa_{B}^{Dh}\kappa_{D}^{f}\cos(\delta_{B}^{Dh}+\delta_{D}^{f}-\gamma)+\Gamma_{\mathrm{mix}}(-x,y,\gamma,f,h),$ (52)
with $\Gamma_{\mathrm{mix}}(x,y,\gamma,f,h)$ defined in Eq. (38). The rate for the CP conjugate mode can be obtained simply replacing $\gamma$ with $-\gamma$ in Eq. (52).
Suppressed (sup) and favoured (fav) decays can be used also to define the following asymmetries
$A^{\text{sup}}(f,h)=\frac{\Gamma(B\to[\overline{f}]_{D}h)-\Gamma(\overline{B}\to[f]_{D}\overline{h})}{\Gamma(B\to[\overline{f}]_{D}h)+\Gamma(\overline{B}\to[f]_{D}\overline{h})},\qquad\qquad A^{\text{fav}}(f,h)=\frac{\Gamma(B\to[f]_{D}h)-\Gamma(\overline{B}\to[\overline{f}]_{D}\overline{h})}{\Gamma(B\to[f]_{D}h)+\Gamma(\overline{B}\to[\overline{f}]_{D}\overline{h})}.$ (53)
In terms of charm parameters, $A^{\text{sup}}(f,h)$ reads
$A^{\text{sup}}(f,h)=\frac{2r_{D}^{f}r_{B}^{Dh}\kappa_{B}^{Dh}\kappa_{D}^{f}\sin\gamma\sin(\delta_{B}^{Dh}+\delta_{D}^{f})+\Gamma_{\mathrm{mix}}^{-}(x,y,\gamma,f,h)}{(r_{B}^{Dh})^{2}+(r_{D}^{f})^{2}+2r_{D}^{f}r_{B}^{Dh}\kappa_{D}^{f}\kappa_{B}^{Dh}\cos\gamma\cos(\delta_{B}^{Dh}+\delta_{D}^{f})+\Gamma_{\mathrm{mix}}^{+}(-x,y,\gamma,f,h)},$ (54)
where $\Gamma_{\mathrm{mix}}^{+}(x,y,\gamma,f,h)$ has been already defined in Eq. (49) for the GLW modes, while $\Gamma_{\mathrm{mix}}^{-}(x,y,\gamma,f,h)$ is given by
$\Gamma_{\mathrm{mix}}^{-}(x,y,\gamma,f,h)=$ $-\alpha r_{B}^{Dh}\kappa_{B}^{Dh}\sin\gamma\bigg{[}y\sin\delta_{B}^{Dh}(1+(r_{D}^{f})^{2})-x\cos\delta_{B}^{Dh}(1-(r_{D}^{f})^{2})\bigg{]}.$ (55)
In the same way, we find for the favoured asymmetry:
$A^{\text{fav}}(f,h)=\frac{2r_{D}^{f}r_{B}^{Dh}\kappa_{B}^{Dh}\kappa_{D}^{f}\sin\gamma\sin(\delta_{B}^{Dh}-\delta_{D}^{f})+\Gamma_{\mathrm{mix}}^{-}(-x,y,\gamma,f,h)}{1+(r_{D}^{f}r_{B}^{Dh})^{2}+2r_{B}^{Dh}\kappa_{B}^{Dh}r_{D}^{f}\kappa_{D}^{f}\cos\gamma\cos(\delta_{B}^{Dh}-\delta_{D}^{f})+\Gamma_{\mathrm{mix}}^{+}(x,y,\gamma,f,h)}.$ (56)
Favoured and suppressed charm modes can be arranged into CP-conserving ratios as
$R^{\text{ADS}}(f,h)=\frac{\Gamma(B\to[\overline{f}]_{D}h)+\Gamma(\overline{B}\to[f]_{D}\overline{h})}{\Gamma(B\to[f]_{D}h)+\Gamma(\overline{B}\to[\overline{f}]_{D}\overline{h})},$ (57)
or using two different channels for beauty decays, $B\to[f]_{D}h$ and $B\to[f]_{D}h^{\prime}$, through the following observables
$R^{\text{sup}}(f,h,h^{\prime})=\frac{\Gamma(B\to[\overline{f}]_{D}h)+\Gamma(\overline{B}\to[f]_{D}\overline{h})}{\Gamma(B\to[\overline{f}]_{D}h^{\prime})+\Gamma(\overline{B}\to[f]_{D}\overline{h^{\prime}})},\qquad\qquad R^{\text{fav}}(f,h,h^{\prime})=\frac{\Gamma(B\to[f]_{D}h)+\Gamma(\overline{B}\to[\overline{f}]_{D}\overline{h})}{\Gamma(B\to[f]_{D}h^{\prime})+\Gamma(\overline{B}\to[\overline{f}]_{D}\overline{h^{\prime}})}.$ (58)
The ADS ratio in Eq. (57) can be expressed in terms of charm parameters as
$R^{\text{ADS}}(f,h)=\frac{(r_{B}^{Dh})^{2}+(r_{D}^{f})^{2}+2r_{D}^{f}r_{B}^{Dh}\kappa_{D}^{f}\kappa_{B}^{Dh}\cos\gamma\cos(\delta_{B}^{Dh}+\delta_{D}^{f})+\Gamma_{\mathrm{mix}}^{+}(-x,y,\gamma,f,h)}{1+(r_{D}^{f}r_{B}^{Dh})^{2}+2r_{B}^{Dh}\kappa_{B}^{Dh}r_{D}^{f}\kappa_{D}^{f}\cos\gamma\cos(\delta_{B}^{Dh}-\delta_{D}^{f})+\Gamma_{\mathrm{mix}}^{+}(x,y,\gamma,f,h)}.$ (59)
The fit equation for $R^{\text{sup}}(f,h,h^{\prime})$ is given by
$R^{\text{sup}}(f,h,h^{\prime})=\frac{\mathcal{B}(B\to D^{0}h)}{\mathcal{B}(B\to D^{0}h^{\prime})}\frac{(r_{B}^{Dh})^{2}+(r_{D}^{f})^{2}+2r_{D}^{f}r_{B}^{Dh}\kappa_{D}^{f}\kappa_{B}^{Dh}\cos\gamma\cos(\delta_{B}^{Dh}+\delta_{D}^{f})+\Gamma_{\mathrm{mix}}^{+}(-x,y,\gamma,f,h)}{(r_{B}^{Dh^{\prime}})^{2}+(r_{D}^{f})^{2}+2r_{D}^{f}r_{B}^{Dh^{\prime}}\kappa_{D}^{f}\kappa_{B}^{Dh^{\prime}}\cos\gamma\cos(\delta_{B}^{Dh^{\prime}}+\delta_{D}^{f})+\Gamma_{\mathrm{mix}}^{+}(-x,y,\gamma,f,h^{\prime})},$ (60)
while for $R^{\text{fav}}(f,h,h^{\prime})$ we have
$R^{\text{fav}}(f,h,h^{\prime})=\frac{\mathcal{B}(B\to D^{0}h)}{\mathcal{B}(B\to D^{0}h^{\prime})}\frac{1+(r_{D}^{f}r_{B}^{Dh})^{2}+2r_{B}^{Dh}\kappa_{B}^{Dh}r_{D}^{f}\kappa_{D}^{f}\cos\gamma\cos(\delta_{B}^{Dh}-\delta_{D}^{f})+\Gamma_{\mathrm{mix}}^{+}(x,y,\gamma,f,h)}{1+(r_{D}^{f}r_{B}^{Dh^{\prime}})^{2}+2r_{B}^{Dh^{\prime}}\kappa_{B}^{Dh^{\prime}}r_{D}^{f}\kappa_{D}^{f}\cos\gamma\cos(\delta_{B}^{Dh^{\prime}}-\delta_{D}^{f})+\Gamma_{\mathrm{mix}}^{+}(x,y,\gamma,f,h^{\prime})}.$ (61)
Appendix C Parameters of the combination
Parameter Range Parameter Range Parameter Range
$\phi_{2}^{M}[\degree]$ $[-11,17]$ ϕ2Γ[°]superscriptsubscriptitalic-$\phi_{2}^{\Gamma}[\degree]$ $[-11,11]$ $F_{+}^{4\pi}$ $[0.5,1]$
$x_{12}\simeq x[\%]$ $[0.05,0.8]$ $y_{12}\simeq y[\%]$ $[0.3,0.9]$ $F_{+}^{\pi\pi\pi^{0}}$ $[0.6,1]$
$a_{D}^{KK}$ $[-0.1,0.1]$ $a_{D}^{\pi\pi}$ $[-0.1,0.1]$ $\Delta Y^{KK}-\Delta Y^{\pi\pi}$ $[-0.1,0.1]$
$(\left\langle\tau\right\rangle^{KK}_{E^{\prime},\pi}+\left\langle\tau\right\rangle^{\pi\pi}_{E^{\prime},\pi})/2$ $[1,3]$ $(\left\langle\tau\right\rangle^{KK}_{E^{\prime},\mu}+\left\langle\tau\right\rangle^{\pi\pi}_{E^{\prime},\mu})/2$ $[1.1,1.3]$ $\left\langle\tau\right\rangle^{KK}_{E^{\prime},\pi}-\left\langle\tau\right\rangle^{\pi\pi}_{E^{\prime},\pi}$ $[0.12,0.15]$
$\left\langle\tau\right\rangle^{KK}_{E^{\prime},\mu}-\left\langle\tau\right\rangle^{\pi\pi}_{E^{\prime},\mu}$ $[-0.01,0.01]$ $\left\langle\tau\right\rangle^{KK}_{E,C_{D^{+}}}$ $[1.46,1.8]$ $\left\langle\tau\right\rangle^{KK}_{E,C_{D_{s}}}$ $[1.6,1.8]$
$\kappa_{D}^{K^{0}_{S}K\pi}$ $[0.3,1]$ $\delta_{D}^{K^{0}_{S}K\pi}[\degree]$ $[-115,115]$ $F_{+}^{KK\pi^{0}}$ $[0.2,1]$
$r_{D}^{K\pi}$ $[0.055,0.065]$ $\delta_{D}^{K\pi}[\degree]$ $[115,258]$ $F_{+}^{KK\pi\pi}$ $[0,1]$
$r_{D}^{K3\pi}$ $[0.048,0.065]$ $\delta_{D}^{K3\pi}[\degree]$ $[29,286]$ $\kappa_{D}^{K3\pi}$ $[0,1]$
$r_{D}^{K\pi\pi^{0}}$ $[0.03,0.06]$ $\delta_{D}^{K\pi\pi^{0}}[\degree]$ $[86,275]$ $\kappa_{D}^{K\pi\pi^{0}}$ $[0.4,1]$
$r_{D}^{K^{0}_{S}K\pi}$ $[0.55,0.7]$
Table 7: Ranges of the uniform priors used for the parameters describing the charm part of the combination. The notation of Sec. III is adopted. Notice that the subscripts $\pi$ and $\mu$ in $\left\langle\tau\right\rangle^{f}_{E^{\prime},\pi}$ and $\left\langle\tau\right\rangle^{f}_{E^{\prime},\mu}$ refer to two different ways of tagging the neutral $D$ mesons. More information are available in ref. Aaij et al. (2019a). The subscripts $C_{D^{+}}$ and $C_{D_{s}}$ in $\left\langle\tau\right\rangle^{f}_{E,D^{+}}$ and $\left\langle\tau\right\rangle^{f}_{E,D_{s}}$ refers to the two strategies adopted in ref. Aaij et al. (2023a) for background subtraction.
Parameter Range Parameter Range Parameter Range Parameter Range
$r_{B}^{DK}$ $[0.07,0.13]$ $r_{B}^{D\pi}$ $[0,0.012]$ $\delta_{B}^{DK}[\degree]$ $[86,172]$ $\delta_{B}^{D\pi}[\degree]$ $[172,360]$
$\mathbf{B^{\pm}}$ $r_{B}^{D^{*}K}$ $[0,0.3]$ $r_{B}^{D^{*}\pi}$ $[0,0.06]$ $\delta_{B}^{D^{*}K}[\degree]$ $[-229,57]$ $\delta_{B}^{D^{*}\pi}[\degree]$ $[-57,303]$
Parameters $r_{B}^{DK^{*\pm}}$ $[0,0.3]$ $\delta_{B}^{DK^{*\pm}}[\degree]$ $[-86,258]$ $\kappa_{B}^{DK^{*\pm}}$ $[0.5,1]$ $r_{B}^{DK\pi\pi}$ $[0,0.2]$
$r_{B}^{D\pi\pi\pi}$ $[0,0.2]$ $\delta_{B}^{DK\pi\pi}[\degree]$ $[-266,94]$ $\delta_{B}^{D\pi\pi\pi}[\degree]$ $[-266,94]$ $\kappa_{B}^{DK\pi\pi}$ $[0,1]$
$\kappa_{B}^{D\pi\pi\pi}$ $[0,1]$ $\frac{\mathcal{B}(B\to DK)}{\mathcal{B}(B\to D\pi)}$ $[0.05,0.1]$
CKM Angles $\gamma[\degree]$ $[40,86]$ $\beta[\degree]$ $[11,29]$ $-2\beta_{s}[\degree]$ $[-17,11]$
$\mathbf{B^{0}}$ $r_{B^{0}}^{DK^{*0}}$ $[0,0.5]$ $\delta_{B^{0}}^{DK^{*0}}[\degree]$ $[115,315]$ $\kappa_{B^{0}}^{DK^{*0}}$ $[0.75,1]$ $r_{B^{0}}^{D\pi}$ $[0,0.4]$
Parameters $\delta_{B^{0}}^{D\pi}[\degree]$ $[-180,180]$
$\mathbf{B^{0}_{s}}$ $r_{B^{0}_{s}}^{D\overline{K^{*0}}}$ $[0,0.5]$ $\delta_{B^{0}_{s}}^{D\overline{K^{*0}}}$ $[-105,255]$ $\kappa_{B^{0}_{s}}^{D\overline{K^{*0}}}$ $[0,1]$ $\delta_{B^{0}_{s}}^{D_{s}K\pi\pi}[\degree]$ $[-115,86]$
Parameters $\kappa_{B^{0}_{s}}^{D_{s}K\pi\pi}$ $[0,1]$ $r_{B^{0}_{s}}^{D_{s}K}$ $[0,1.2]$ $\delta_{B^{0}_{s}}^{D_{s}K}[\degree]$ $[-143,143]$ $r_{B^{0}_{s}}^{D_{s}K\pi\pi}$ $[0,1.2]$
Table 8: Ranges of the uniform priors employed for the parameters describing the beauty observables, divided by the species of the $B$ mesons involved. The notation of Sec. IV is adopted.
We report the intervals defining the uniform priors adopted for the charm parameters in Tab. 7, while the beauty parameters are contained in Tab. 8, divided by types of $B$ mesons involved in the observables. The intervals are chosen according to the definitions of the parameters and to speed up the convergence of the Markov chains, after having checked that no solutions are lost.
Appendix D Fit results
Figure 6: Two-dimensional contours of the charm and beauty decay parameters describing the most precise observables of the combination and the direct CP asymmetries in $D^{0}\to K^{+}K^{-}$ and $D^{0}\to\pi^{+}\pi^{-}$ decays. Darker (lighter) regions correspond to $68.3\%$ ($95.4\%$) probability.
$\gamma$ $x_{12}$ $y_{12}$ $\phi_{2}^{M}$ $\phi_{2}^{\Gamma}$ $\phi_{12}$ $r_{B}^{DK}$ $\delta_{B}^{DK}$ $r_{B}^{D\pi}$ $\delta_{B}^{D\pi}$ $r_{D}^{K\pi}$ $\delta_{D}^{K\pi}$ $\phi_{2}$ $|q/p|-1$ $a_{D}^{KK}$ $a_{D}^{\pi\pi}$
$\gamma$ $100$- $1$- -- -- $43$ $50$ $14$ $24$- $4$- -- -- --
$x_{12}$ $100$ $4$ $-12$ $1$ $-10$ $1$ $2$ $-2$ $-1$ $8$ $2$ $5$ $-10$ $1$-
$y_{12}$ $100$ $-1$ $-6$ $2$ $-10$ $-20$ $4$ $-2$ $-14$ $42$ $6$ $2$- --
$\phi_{2}^{M}$ $100$ $-4$ $81$- $1$- -- $-3$ $-3$ $-44$ $81$ $31$ $26$
$\phi_{2}^{\Gamma}$ $100$ $-62$ $1$ $3$- $1$- $-5$ $-87$ $-62$ $-1$ $-1$
$\phi_{12}$ $100$- $-1$- -- $-3$ $1$ $17$ $100$ $24$ $21$
$r_{B}^{DK}$ $100$ $33$ $4$ $11$ $-9$ $-21$ $-1$- -- --
$\delta_{B}^{DK}$ $100$ $8$ $23$ $-14$ $-51$ $-3$ $-1$- --
$r_{B}^{D\pi}$ $100$ $68$ $13$- -- -- -- --
$\delta_{B}^{D\pi}$ $100$ $11$ $-18$- -- -- --
$r_{D}^{K\pi}$ $100$ $28$ $1$ $-3$- --
$\delta_{D}^{K\pi}$ $100$ $6$ $1$- --
$\phi_{2}$ $100$ $17$ $-14$ $-12$
$|q/p|-1$ $100$ $24$ $21$
$a_{D}^{KK}$ $100$ $67$
$a_{D}^{\pi\pi}$ $100$
Table 9: Correlation matrix for the most significant quantities in percent. Values smaller than $1\%$ are not reported. Large correlations between $\phi_{2}^{M,\Gamma}$, $\phi_{12}$, $|q/p|-1$ and $\phi_{2}$ are due to the fact that they are not independent parameters, as shown in Eqs. (7) and (13).
We report the results of the fit obtained by combining all the observables.
The correlation matrix of the most physically relevant parameters is reported in Tab. 9.
Tabs. 10 and 11 present the probability intervals found for the charm and beauty parameters, respectively. In Fig. 6, we report two-dimensional contours of the ratio of decay amplitudes and strong phases describing the most precise modes used in the combination: $B^{\pm}\to D\pi^{\pm}(K^{\pm})$ and $D\to K^{\pm}\pi^{\mp}$, together with the direct CP asymmetries between the charm decay amplitudes of the modes $K^{+}K^{-}$ and $\pi^{+}\pi^{-}$.
Parameter Value Unc. $\mathbf{95.4\%}$ Prob. Parameter Value Unc. $\mathbf{95.4\%}$ Prob.
$\phi_{2}^{M}[\degree]$ $2.4$$\pm 1.6$ $[-0.7,5.7]$ ϕ2Γ[°]superscriptsubscriptitalic-$\phi_{2}^{\Gamma}[\degree]$ $2.6$$\pm 1.2$ $[0.2,5.0]$
$F_{+}^{4\pi}$ $0.745$$\pm 0.010$ $[0.725,0.764]$ $x_{12}\simeq x[\%]$ $0.402$$\pm 0.034$ $[0.335,0.470]$
$y_{12}\simeq y[\%]$ $0.629$$\pm 0.021$ $[0.587,0.672]$ $F_{+}^{\pi\pi\pi^{0}}$ $0.973$$\pm 0.015$ $[0.944,1.000]$
$a_{D}^{KK}[\%]$ $0.085$$\pm 0.063$ $[-0.041,0.211]$ $a_{D}^{\pi\pi}[\%]$ $0.301$$\pm 0.070$ $[0.161,0.443]$
$F_{+}^{KK\pi^{0}}$ $0.738$$\pm 0.055$ $[0.629,0.847]$ $r_{D}^{K\pi}[\%]$ $5.845$$\pm 0.017$ $[5.812,5.878]$
$\delta_{D}^{K\pi}[\degree]$ $191.1$$\pm 2.9$ $[185.3,196.8]$ $F_{+}^{KK\pi\pi}$ $0.830$$\pm 0.071$ $[0.691,0.969]$
$r_{D}^{K3\pi}[\%]$ $5.564$$\pm 0.058$ $[5.448,5.682]$ $\delta_{D}^{K3\pi}[\degree]$ $157$$\pm 15$ $[127,187]$
$\kappa_{D}^{K3\pi}$ $0.478$$\pm 0.064$ $[0.346,0.606]$ $r_{D}^{K\pi\pi^{0}}[\%]$ $4.43$$\pm 0.11$ $[4.22,4.64]$
$\delta_{D}^{K\pi\pi^{0}}[\degree]$ $203.3$$\pm 9.0$ $[185.1,221.2]$ $\kappa_{D}^{K\pi\pi^{0}}$ $0.794$$\pm 0.040$ $[0.714,0.873]$
$r_{D}^{K^{0}_{S}K\pi}$ $0.6196$$\pm 0.0053$ $[0.6090,0.6304]$ $\delta_{D}^{K^{0}_{S}K\pi}[\degree]$ $13$$\pm 14$ $[-16,40]$
$\kappa_{D}^{K^{0}_{S}K\pi}$ $0.821$$\pm 0.094$ $[0.647,0.996]$ $|q/p|-1[\%]$ $-0.1$$\pm 1.6$ $[-3.3,3.1]$
ϕ2[°]subscriptitalic-$\phi_{2}[\degree]$ $-2.54$$\pm 0.94$ $[-4.44,-0.66]$ ϕ12[°]subscriptitalic-$\phi_{12}[\degree]$ $-0.2$$\pm 2.0$ $[-4.1,4.0]$
Table 10: Results for the charm part of the combination. For each of the parameters, we have reported the central value (“Value”) and the smallest intervals containing the $68.3\%$ (“Unc.”) and $95.4\%$ probabilities.
Parameter Value Unc. $\mathbf{95.4\%}$ Prob. Parameter Value Unc. $\mathbf{95.4\%}$ Prob.
$r_{B}^{DK}[\%]$ $9.99$$\pm 0.23$ $[9.53,10.46]$ $r_{B}^{D\pi}[10^{-3}]$ $5.05$$\pm 0.72$ $[3.78,6.79]$
$\mathbf{B^{\pm}}$ $\delta_{B}^{DK}[\degree]$ $128.7$$\pm 2.9$ $[122.6,134.2]$ $\delta_{B}^{D\pi}[\degree]$ $297$$\pm 12$ $[270,319]$
Parameters $r_{B}^{D^{*}K}$ $0.109$$\pm 0.011$ $[0.087,0.129]$ $r_{B}^{D^{*}\pi}[\%]$ $0.59$$\pm 0.34$ $[0,1.3]$
$\delta_{B}^{D^{*}K}[\degree]$ $-44.0$$\pm 6.7$ $[-58.6,-31.6]$ $\delta_{B}^{D^{*}\pi}[\degree]$ $51$$\pm 41$ $[-7,161]$
$r_{B}^{DK^{*\pm}}$ $0.098$$\pm 0.009$ $[0.078,0.115]$ $\delta_{B}^{DK^{*\pm}}[\degree]$ $49$$\pm 12$ $[27,79]$
$\kappa_{B}^{DK^{*\pm}}$ $0.957$$\pm 0.040$ $[0.856,1.000]$ $\frac{\mathcal{B}(B\to DK)}{\mathcal{B}(B\to D\pi)}[\%]$ $7.79$$\pm 0.37$ $[7.05,8.54]$
CKM Angles $\gamma[\degree]$ $68.4$$\pm 2.6$ $[63.2,73.5]$ $\beta[\degree]$ $22.33$$\pm 0.57$ $[21.19,23.48]$
$-2\beta_{s}[\degree]$ $-2.7$$\pm 1.1$ $[-4.9,-0.6]$
$\mathbf{B^{0}}$ $r_{B^{0}}^{DK^{*0}}$ $0.240$$\pm 0.012$ $[0.215,0.265]$ $\delta_{B^{0}}^{DK^{*0}}[\degree]$ $192.9$$\pm 4.9$ $[183.1,203.1]$
Parameters $\kappa_{B^{0}}^{DK^{*0}}$ $0.957$$\pm 0.023$ $[0.914,1.000]$ $r_{B^{0}}^{D\pi}$ $0.027$$\pm 0.012$ $[0.002,0.050]$
$\delta_{B^{0}}^{D\pi}[\degree]$ $25$$\pm 30$ $[-51,80]$
$r_{B^{0}_{s}}^{D_{s}K}$ $0.328$$\pm 0.035$ $[0.259,0.399]$ $\delta_{B^{0}_{s}}^{D_{s}K}[\degree]$ $-11.5$$\pm 6.0$ $[-23.7,0.6]$
$\mathbf{B^{0}_{s}}$ $r_{B^{0}_{s}}^{D_{s}K\pi\pi}$ $0.449$$\pm 0.078$ $[0.294,0.604]$ $\delta_{B^{0}_{s}}^{D_{s}K\pi\pi}[\degree]$ $-16$$\pm 13$ $[-41,10]$
Parameters $\kappa_{B^{0}_{s}}^{D_{s}K\pi\pi}$ $0.74$$\pm 0.14$ $[0.50,1.00]$ $r_{B^{0}_{s}}^{D\overline{K^{*0}}}$ $0.060$$\pm 0.030$ $[0,0.101]$
$\delta_{B^{0}_{s}}^{D\overline{K^{*0}}}[\degree]$ $80$$\pm 72$ $[-62,239]$ $\kappa_{B^{0}_{s}}^{D\overline{K^{*0}}}$ $0.3$$\pm 0.3$ $[0,0.9]$
Table 11: Results for the beauty part of the combination. For each of the parameters, we have reported the central value (“Value”) and the smallest intervals containing the $68.3\%$ (“Unc.”) and $95.4\%$ probabilities.